### 06 Mar 2026 v2

# “When to Hand Off, When to Work Together”: Expanding Human-Agent Co-Creative Collaboration through Concurrent Interaction

Abstract.
Human collaborators coordinate dynamically through process visibility and workspace awareness, yet AI agents typically either provide only final outputs or expose read-only execution processes (e.g., planning, reasoning) without interpreting concurrent user actions on shared artifacts. Building on mixed-initiative interaction principles, we explore whether agents can achieve collaborative context awareness—interpreting concurrent user actions on shared artifacts and adapting in real-time. Study 1 (N=10 professional designers) revealed that process visibility enabled reasoning about agent actions but exposed conflicts when agents could not distinguish feedback from independent work. We developed Cleo, which interprets collaborative intent and adapts in real-time. Study 2 (N=10, two-day with stimulated recall interviews) analyzed 214 turns, identifying five action patterns, six triggers, and four enabling factors explaining when designers choose delegation (70.1%), direction (28.5%), or concurrent work (31.8%). We present a decision model with six interaction loops, design implications, and an annotated dataset.
Human-Agent Collaboration, Interaction Pattern Analysis, Concurrent Human-Agent Interaction, Generative AI, Co-Creative Design Agent
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Interactive systems and tools
1. Introduction
In human–human collaborative creative work, as commonly observed in shared tools such as Figma [^1] https://www.figma.com or Google Slides [^2] https://docs.google.com/presentation, visibility into a collaborator’s actions and awareness of who is working on what are fundamental for effective coordination (Dourish and Bellotti, 1992). Collaborators observe each other’s ongoing work, anticipate next steps, intervene at opportune moments, and fluidly divide tasks based on emerging priorities and expertise (Tang, 1991). This dynamic coordination relies on process transparency—the ability to see how work unfolds, not just what is produced—and mutual awareness of presence and focus within the shared workspace (Gutwin and Greenberg, 2002; Gutwin et al., 1996). These elements enable richer collaboration patterns beyond sequential turn-taking, including concurrent work, interleaved workflows, and real-time adaptation to each other’s contributions (Tang, 1991).
AI agents increasingly perform tasks step-by-step using human-interpretable operations (e.g., checking files, clicking buttons), making their processes observable across domains like programming (Sapkota et al., 2025), web automation (Jiang and Xia, 2025), and research (Gottweis et al., 2025). In creative domains, however, agents typically provide only final outputs rather than exposing execution processes (Make, ; Stitch, ; Lovable, ), limiting them to assistant roles for specific subtasks (You et al., 2025; Zhou et al., 2024). Yet process observability alone is insufficient. The conceptual challenge of flexibly shifting between delegation and direct control based on context has deep roots in mixed-initiative interaction (Horvitz, 1999) and adjustable autonomy in robotics (Pellegrinelli et al., 2016). What remains unexplored is whether agents in creative contexts can go beyond read-only process transparency to achieve collaborative context awareness—interpreting concurrent user direct manipulation on shared artifacts and adapting execution in real-time, enabling the fluid, interleaved collaboration patterns seen in human–human creative work.
We investigate this question through two design probes in Figma. Our first probe established process transparency through voice interaction, visual tool operations, and workspace awareness (agent cursor presence). Study 1 (N=10 professional designers) revealed that while step-by-step execution visibility enabled anticipation and timely intervention, designers expected mid-execution concurrent collaboration—wanting to interleave their work with the agent’s ongoing execution. The agent, however, could not distinguish user feedback from independent work, creating conflicts that revealed a critical gap: the absence of collaborative context awareness.
This finding motivated Cleo (Collaborative Linked Executive Operator), which tracks concurrent user actions, interprets collaborative intent (feedback vs. independent work), and adapts in real-time. Study 2 (N=10, two-day exploratory study with stimulated recall interviews (Dempsey, 2010)) analyzed 214 turn-level interactions to understand when and why designers choose different collaboration modes. Through qualitative coding, we identified five action patterns (hands-off, observational, concurrent, directive, terminating in Figure 7), six triggers prompting intervention (e.g., idea spark, task mismatch in Figure 8), and four enabling factors determining choices (e.g., mental model, task priority in Figure 9). Concurrent work appeared in 31.8% of turns, full delegation in 70.1%, and directive intervention in 28.5%, with dynamic shifts based on evolving priorities.
We synthesized these findings into a decision model (Figure 10) that characterizes how designers dynamically navigate among five actions: delegation, observation, direction, concurrent work, and termination. The model identifies six interaction loops driven by triggers, enabling factors, and evolving priorities. Drawing from the dynamics revealed by these loops, we derive design implications for adaptive collaborative agents that support seamless mode transitions and context-aware adaptation. We also release our dataset of 214 interaction turns with logged actions, triggers, and designer rationales to support future research in human-agent co-creative collaboration.
Our contributions include:
  • A characterization of concurrent human-agent interaction with five user action patterns enabling fluid collaboration during agent execution.
  • A decision model with six interaction loops explaining how users navigate between delegation, direction, and concurrent work based on six triggers and four enabling factors.
  • Design implications for adaptive collaborative agents supporting seamless mode transitions and context-aware adaptation.
  • A dataset of 214 human-agent co-creative interaction turns with logged actions.
2. Related Work
2.1. Dynamics of Flexible Collaboration: From Turn-taking to Synchrony
Human collaboration is a dynamic spectrum shifting between different interaction modes (Tang, 1991). While turn-taking facilitates coordination in complex tasks by minimizing cognitive load and conflicts through sequential actions (Kirschner et al., 2009; Sacks et al., 1974), creative domains often demand synchronous collaboration (Nam and Sakong, 2009). This is because creative work is rarely a linear process of executing predefined steps; rather, it involves the co-evolution of problem and solution, where design ideas are refined through immediate visual feedback and spontaneous mutual inspiration (Maher and Poon, 1996; Ishii et al., 1993). In such contexts, waiting for a turn can break the creative flow and delay the rapid synthesis of ideas. Therefore, partners often bump in for immediate adjustments or perform simultaneous edits to interleave workflows, ensuring that the evolving design remains a shared mental model (Dourish and Bellotti, 1992; Tang, 1991).
Modern tools like Figma and Google Slides support this fluidity by fostering workspace awareness—the real-time understanding of a collaborator’s presence and intentions. This is enabled through visual cues like telepointers (moving cursors) and object highlighting, allowing users to anticipate moves and intervene without explicit communication (Gutwin and Greenberg, 2002; Gutwin et al., 1996). However, most co-creative AI agents remain confined to a rigid wait-your-turn paradigm. Unlike the parallel dynamics mastered in human-human work, human-AI interaction is still largely limited to a prompt-and-output cycle, lacking the mutual awareness essential for true creative partnership.
$2.$ Paradigm Shifts in Human-Agent Collaboration
AI is evolving from a command-based tool into a proactive “co-worker” that actively participates within shared workspaces (Horvitz, 1999). This shift has been driven by increased AI capabilities and frameworks that enable models to leverage diverse tools (e.g., code, web search, etc.) (Schick et al., 2023) and interleave reasoning with actions across extended time horizons (Yao et al., 2023). These capabilities have enabled diverse agents for programming (by Anthropic, ; Jimenez et al., 2023; Sapkota et al., 2025), web automation (AI, ; Deng et al., 2023; Jiang and Xia, 2025), visual design (Make, ), research (Gottweis et al., 2025), data analysis (Hong et al., 2025) etc. As these agents expose their reasoning and intermediate actions within their generation, this naturally affords greater interpretability by allowing users to monitor the agent behavior in real time (Amershi et al., 2019; Prasongpongchai et al., 2025).
However, despite this increased visibility, most collaboration between human users and AI agents remains rooted in turn-taking. Even when users can interpret what actions the agent is taking on which artifacts, they can rarely act simultaneously on those artifacts. This limits users’ ability to correct agents mid-execution (Amershi et al., 2019) to prevent misalignments as the task progresses. In this work, we explore the potential of concurrent human-agent collaboration not as a replacement for turn-taking, but as an additional mode for collaboration—enabling agents to move beyond delegated executors to responsive partners capable of accommodating spontaneous interventions within a shared workspace.
2.3. Design Agents: From Workflow Automation to Process-Visible Collaboration
Existing agents for supporting the design process can be categorized into three distinct approaches based on how they structure their tasks and interact with users. The first category is subtask design agents, which support specific stages within an existing design workflow by automating specific and tedious subtasks, such as adding the details (Lawton et al., 2023; Karimi et al., 2020; Zhou et al., 2024) or generating diverse alternatives (Make, ). In this model, the overall workflow remains human-driven, while the agent functioning as an assistive role. The second category is process-structured agents, which operate within a predefined step-by-step design activities through node-based workflow (Hao et al., 2025; Wang et al., 2025; Ding et al., 2023; You et al., 2025). These systems expose intermediate process steps within the whole workflow and support guided co-creation between agent and human, but the collaboration is constrained by predefined logical nodes. The third type is the output-centric agent, which exposes an internal reasoning process for generating a final output to users in text, such as their execution planning (Stitch, ; Lovart, ; Khurana et al., 2025).
Despite these differences in task structuring, these agents share a common interaction characteristic that the agent’s execution process is hidden from user’s workspace. Users interact with agents only through requests and final outputs rather than through ongoing execution steps. As a result, users are required to wait for final results without seeing the work in progress. While agents in other domains (e.g., programming) increasingly expose execution steps through direct tool operations, design tools have yet to explore how “live” tool-use—such as visible cursor movements and object manipulation—could reshape human-agent collaboration. This suggests an open question of whether process visibility could enable more diverse collaboration modes beyond simple turn-taking.
3. Study 1: Initial Technology Probe Study
To understand how agents’ workflow sharing via human-understandable tool calls affects human-agent collaboration in design, we implemented an initial design probe and conducted an exploratory study with 10 professional designers.
3.1. First Probe System
3.1.1. Three Basic Components for Collaborative Agent Design
We ground the system’s architecture in three fundamental components from human-human collaboration for collaborative agent design: (1) Interaction Modality for peer-like communication, (2) Process Transparency for observable agent work, and (3) Presence Awareness for mutual attention visualization.
Interaction Modality: Voice Interaction with Flexible Object Referencing
The first probe system uses voice-based interaction as the primary channel, enabling natural turn-taking and real-time intention sharing (Hutchins, 1995; Reicherts et al., 2022). Voice creates a sense of horizontal partnership rather than task delegation (Bickmore and Picard, 2005). Furthermore, users can communicate with the system through direct object referencing on the canvas, which enables fluid deictic communication (e.g., “make this larger” while clicking).
Process Transparency: Sharing Implementation Workflow
Unlike general AI design tools that only share final outputs, the system makes its design process visible in real-time through the execution of human design tools (e.g., creating a rectangle to make a button). As the agent generates UI components step by step, users observe elements appearing, positioning, and styling incrementally—ensuring “public availability of actions and artifacts” (Robertson, 2002). This progressive visualization reveals not just what the agent produces, but how it is progressing, establishing common ground around both goals and processes (Cila, 2022; Hill et al., 1992) between the user and the agent.
Presence Awareness: Visualizing Attention in the Shared Workspace
The system displays real-time indicators of agent’s current focus area and working state, following Gutwin and Greenberg’s workspace awareness framework (Gutwin and Greenberg, 2002). The agent character’s location reveals which canvas region the agent is attending to, enabling users to anticipate the next execution area and coordinate accordingly (Dourish and Bellotti, 1992; Fussell et al., 2000). This peripheral awareness creates mutual presence beyond simple loading indicators.
3.1.2. System Design
Figure 1. Interaction flow of the first probe system. (a) The user clicks the record button to deliver a voice message (b) while selecting canvas elements for context. The transcribed message with selection changes is sent to the agent as input. The agent generates an action plan and iterates through cycles of (1) reason, (2) act, and (3) feedback. (d-f) Each iteration’s progress is reflected immediately on the canvas as tools execute during the act stage, with the agent character floating near modified elements and displaying interim messages.
This figure illustrates the interaction flow of the first probe system through an example scenario in which a user requests color variations for a selected section. The user delivers a voice request while selecting relevant canvas elements, allowing the system to capture both the verbal intent and the visual context. Based on this input, the agent generates an action plan and iteratively progresses through cycles of reasoning, tool execution, and feedback evaluation. During each iteration, the agent executes design tools directly on the canvas and provides interim messages, enabling users to observe incremental changes and intervene if needed before the final output is produced.
Interface and Interaction Scenario
The first probe system is implemented as a Figma plugin. Users interact with the system through both the plugin interface and an agent character that appears on the canvas. To provide an input to the agent, users click the record button in the plugin interface to send voice messages (Figure 1A). The system continuously tracks voice input and selection changes on the canvas until the user clicks the stop button. Users can click any element, including reference images or specific frames, to provide accurate context. As illustrated in Figure 1B, when a user says “show color variation of this section” while selecting an element, the transcribed message, along with timestamp and selected element metadata, is sent to the agent as input. Upon receiving user input, the agent responds with a sequence of Figma operations and text messages. The agent completes its response through multiple iterations, executing actions and providing messages incrementally (Figure 1C). For example, rather than displaying style variations all at once, the agent first duplicates the original element (Figure 1D), then applies style changes sequentially (Figure 1E-F), revealing progress step-by-step. At each step, the agent character animates around the modified elements with interim messages, helping users easily track the agent’s actions.
Figure 2. Agent structure of the first probe system. Upon receiving user input, the Plan Module generates an action plan. The ReAct Agent then iterates through a cycle of reasoning with tool selection, execution, summary, and feedback evaluation until the plan is fulfilled or the maximum of 10 iterations is reached. The Message Module generates the final response, reflecting both the agent’s actions and the user’s original message, and delivers it to the user.
This figure presents the internal architecture of the first probe system and the control flow of the ReAct-based agent. User input is first combined with the chat history and canvas context, which are used by the Plan Module to generate an action plan. The ReAct Agent then executes the plan through an iterative loop that includes reasoning with tool selection, tool execution on the canvas, summarization of tool call results, and feedback evaluation. Interim messages and updated canvas context are produced during execution to reflect the agent’s ongoing progress. This loop continues until the plan is fulfilled or a maximum of ten iterations is reached, after which the Message Module generates a final response that integrates the agent’s actions with the original user request.
Agent Design and Pipeline
The first probe agent is designed based on the ReAct agent structure (Yao et al., 2023). As illustrated in Figure 1C, it operates through a three-step iteration cycle: (1) reason, (2) act, and (3) feedback. Referring to Figure 2, the cycle begins with the main reasoning stage, which determines whether design actions are required. If the agent deems actions unnecessary, it terminates the cycle and provides a final response through the Message Module. Otherwise, it selects tools from the 38 Figma operations and generates an interim progress message. The selected tools are immediately executed on the canvas alongside the agent character and the interim message.
After each iteration cycle, the summary module generates a text description of what occurred based on the tool execution results. These summaries are retained in context and stored with the chat history. The feedback module then captures the current canvas state and compares it with the previous state before tool execution to assess whether the plan has been fulfilled, and generates a feedback message describing the next steps if needed. If the plan is complete, the agent terminates the loop and proceeds to the message module. Otherwise, it returns to the initial stage of the iteration cycle to begin a new iteration. To prevent infinite loops and ensure usability, we limit the maximum number of iteration cycles to 10.
We utilize Anthropic’s Claude Sonnet 4.5  [^3] https://www.anthropic.com/claude/sonnet as the main reasoning model and Claude Haiku 4.5  [^4] https://www.anthropic.com/claude/haiku for all modules including plan, summary, feedback, and message generation. To enhance reasoning performance, we provide the main model with guidance instructions for Figma, including recommended rules and patterns for UI design. Detailed instructions for each module and the complete set of 38 Figma operation tools can be found in the Appendix. We designed the prompts (e.g., Figma design guideline) by referring Talk To Figma plugin  [^5] https://github.com/grab/cursor-talk-to-figma-mcp.
3.2. Study Design
Our first study aims to investigate two key questions: (1) How does process transparency—observing the agent’s step-by-step workflow on the canvas—affect designers’ collaboration experience and enable a new type of collaboration? (2) What design considerations are needed to support better human-agent collaboration in co-creative design tasks?
To answer these questions, we recruited 10 professional UI/UX designers and had them work with our design probe on two types of design tasks, each representing a different level of task granularity. Through task observations and semi-structured interviews, we investigated how process transparency affects designers’ understanding, collaboration strategies, and expectations for agent behavior.
3.2.1. Participants
We recruited 10 UI/UX designers (D1-D10, 3 male, 7 female, ages 24-36, mean=28.30) with professional design experience. All participants had at least 1 year of professional experience or had completed at least three web or UI/UX design projects. We screened these participants for high Figma proficiency (e.g., daily users) and prior experience with diverse design agents (e.g., Figma Make (Make, ) or Lovart (Lovart, )). Participants were compensated 40,000 KRW for approximately 1.5 hours of participation.
3.2.2. Procedure
After introducing different types of design agents (e.g., OptiMuse (Zhou et al., 2024) and Figma Make (Make, )) and providing a brief tutorial on the design probe’s three interaction components (voice instruction, process observation, and presence visualization), participants completed two design tasks of varying scope: (1) a fine-grained task, which involved designing a single webpage section (e.g., hero or project gallery section) with given design materials (15 minutes), and (2) a coarse-grained task, which involved creating an entire webpage (e.g., startup landing page) including multiple sections (25 minutes). These tasks were designed to observe how collaboration methods differ between single-section and multi-section work. We concluded with a 30-minute semi-structured interview exploring how observing the agent’s workflow affected their collaboration strategies, their experience instructing the agent, and desired improvements for the current design probe.
$3.$ Findings
3.3.1. Process Transparency: Real-Time Understanding and Intervention Opportunities
Observing the agent’s step-by-step execution fundamentally transformed how designers experienced collaboration. All designers reported a strong sense of working with the agent rather than delegating to it. Unlike traditional agents, where users assign tasks and wait for final outputs, process transparency enabled simultaneous work: participants understood what the agent was doing and planned their own tasks accordingly, maintaining active engagement throughout execution.
This transparency relieved the “loss of agency” problem common in existing agents (Feng et al., 2024). Participants contrasted the first probe system with general design agents, where agents take over the workspace, leaving users waiting passively and fearing disruption if they intervene. With those agents, users must read lengthy reasoning texts to understand its plans—a cognitively demanding process. In contrast, seeing intermediate visual actions on the canvas made the agent’s intentions immediately understandable with minimal effort. D4 commented “I don’t need to read paragraphs—I just see it creating a button, and I know what’s happening.” This visual feedback preserved user agency by avoiding both over-dependence on opaque automation and wholesale rejection of agent outputs.
Most critically, step-by-step observation enabled early evaluation and proactive intervention. Participants could anticipate the agent’s next steps, identify the agent’s misinterpretation on the given task (D1, D3, and D5-D7), recognize their own unclear instructions (D1-D2, D4, and D8), and assess the agent’s overall workflow direction (D2, D6, and D8-D10)—all before the agent finished. Importantly, this understanding enabled participants to plan their own parallel work: while the agent executed, participants gathered reference images, evaluated overall design directions, or prepared content for subsequent sections, rather than passively waiting. Specifically, D7 explained: “I saw it working on the hero section with a certain direction, so I started looking for references related to that direction for the next section instead of just watching.” This predictability helped participants manage workflows more effectively, deciding when to intervene, when to prepare subsequent tasks, and when to let the agent continue autonomously. Process transparency thus shifted collaboration from reactive correction of final outputs to proactive coordination during execution.
3.3.2. Emerging Expectations: Flexible Concurrent Collaboration and Context Awareness
Agent’s process transparency fundamentally changed designers’ expectations: they wanted to engage directly with the agent’s work-in-progress through concurrent intervention rather than waiting for completion. For example, participants wanted to appropriate intermediate outputs for their own designs, assist the agent by working on related elements simultaneously, or demonstrate desired changes through direct editing. Most of the participants (D1-D3 and D6-D10) explained that observing the process made them more aware of how they wanted things done, creating natural moments to intervene and collaborate. D3 noted: “I could see exactly where it needed my input—I wanted to just jump in and show it rather than wait and explain later.”
However, the first probe’s inability to distinguish user actions from agent actions created conflicts. Because the agent tracked overall canvas changes without separating user modifications, it sometimes reverted or misinterpreted concurrent user work. Participants who experienced this avoided intervening during execution, relegating their collaboration ideas to post-task interviews. Yet even participants who avoided intervention reported that their reluctance to interrupt agents stems from prior experiences with other agent systems (D2, D4-D7, D10)—fearing that mid-execution changes would derail the entire agent’s workflow. Process transparency challenged this assumption; designers recognized opportune moments for intervention but lacked system support.
Participants envisioned real-time interleaved workflows where users could intervene and interleave their work with the agent’s ongoing execution: appropriating agent-created elements into their own designs while the agent continued working, handing off their work-in-progress for the agent to build upon mid-execution, or collaboratively editing the same artifact simultaneously. Critically, they expected agents to understand why they intervened—whether corrections indicated misunderstanding, appropriations showed approval, or demonstrations communicated desired style (D1, D3-D5, D7-D8). This requires collaborative context awareness, which means distinguishing feedback on agent work from independent user work and recognizing when users want the agent to adapt versus continue autonomously. Process transparency thus revealed not just new collaboration opportunities, but the need for context-aware agents that support—rather than conflict with—real-time interleaved human actions.
4. Study 2: Two-Day Technology Probe Study with Stimulated Recall Interviews
In this section, we describe an exploratory study that investigates when designers delegate hands-off tasks to agents, work collaboratively alongside them, or provide only verbal feedback during co-creative collaboration. Through a two-day technology probe study (Hutchinson et al., 2003) incorporating stimulated recall interviews (Dempsey, 2010; Calderhead, 1981) with creative professionals having over two years of experience, we sought to understand how designers navigate various action patterns in collaborative contexts with the agent, where natural, simultaneous work with agents is possible. This section first describes the refined design probe used in this study, followed by the study design, data analysis methods, and findings.
4.1. Second Probe System: Cleo (Collaborative Linked Executive Operator)
4.1.1. Understanding Collaborative Context for Flexible Human-Agent Co-Creation
While our first design probe includes three components (voice modality, process transparency, and presence awareness), supporting real-time interleaved workflows requires agents to understand and adapt to diverse human interventions that occur during the agent’s execution. Based on Study 1 findings, users expect to interleave their actions with the agent’s ongoing execution rather than waiting for task completion when they can observe and understand the agent execution process. The key challenge is collaborative context awareness—agents must distinguish whether concurrent user actions represent feedback on the agent’s work requiring adaptation, the user’s independent parallel work that should be preserved, or collaborative co-editing that should be incorporated into the agent’s ongoing execution. This requires that the agent should understand which elements users modify, whether they build upon or diverge from the agent’s work, and how users’ demonstrations relate to the ongoing task.
Therefore, we developed a second design probe, Cleo (Collaborative Linked Executive Operator), which extends the first probe by incorporating collaborative context awareness as its core principle. This enables the agent to recognize the user’s direct manipulation of the agent’s execution process and adapt its behavior accordingly while maintaining coherent shared progress.
4.1.2. Cleo Design and Improvements from Design Probe 1
Figure 3. Flexible co-creation scenario with Cleo. (a) The user invokes the agent by calling its name. Cleo allows users to (b) abort the agent’s operation, (c) work concurrently with the agent, or (d) intervene directly with additional instructions at any time.
This figure illustrates a flexible co-creation scenario enabled by Cleo, highlighting how users can dynamically shift their level of involvement during an ongoing task. After invoking the agent to create a portfolio header, the user observes the agent’s initial output on the canvas. In subsequent iterations, the user may abort the agent’s operation to work independently, make concurrent edits directly on the canvas while the agent is active, or provide additional verbal instructions to redirect the agent’s behavior. The sequence demonstrates how Cleo supports fluid transitions between hands-off delegation, concurrent collaboration, and direct intervention within a single task.
Interaction Improvements
Building on the first probe system, Cleo enables more dynamic and flexible collaboration as illustrated in Figure 3. Users can invoke the agent by saying “Cleo” without clicking the microphone button. The system then detects voice input and tracks canvas selections until several seconds of silence occur.
Cleo supports three types of interactions. First, users can abort the agent’s operation at any time via the termination button in the plugin interface (Figure 3B). Second, users can work concurrently with the agent, even on the same elements (Figure 3C). The agent tracks and prioritizes user operations over its own actions, updating its plan to align with user modifications. For instance, as shown in Figure 3C(2)-(3), when a user changes the title color to blue while the agent works on the menu, the agent incorporates blue as a key color in subsequent actions, such as adding blue decorative lines. Third, users can provide additional voice instructions at any time by calling the agent’s name (Figure 3D), which the agent incorporates in the next iteration.
Technical Improvements
To support these additional interactions, we updated the agent pipeline as shown in Figure 4. Instead of the original Summary Module, which only considers the agent’s tool calls and results, we introduced three modules to incorporate user actions into the agent’s workflow planning: Attribution Change Module, User Change Detection Module, and Plan Update Module.
When the agent receives user input, the User Change Detection Module analyzes user actions that occurred between the previous agent response and the current canvas state, capturing changes made while the agent was inactive. This allows the agent to maintain awareness of the overall workflow. The Attribution Change Module distinguishes between agent operations and user modifications by analyzing the canvas context after tool execution. It produces two separate summaries—one for agent actions and one for user actions during the iteration. These summaries are passed to the Plan Update Module, which updates the current plan only when user modifications conflict with the existing plan or indicate new directions, before initiating the next iteration cycle.
Figure 4. Agent structure of Cleo. Three modules have been updated from the first probe pipeline: User Change Detection Module, Attribution Change Module, and Plan Update Module.
This figure presents the updated agent architecture of Cleo, highlighting three additional modules introduced beyond the first probe pipeline to support user interventions during agent execution. Before planning, the User Change Detection Module analyzes user actions that occurred while the agent was inactive and incorporates them into the chat history and canvas context. During execution, the Attribution Change Module distinguishes between agent-generated changes and user modifications by processing tool call results together with the current canvas state. Based on this attribution, the Plan Update Module selectively revises the agent’s plan when user actions conflict with the current plan or indicate new directions. These updates are integrated into the existing ReAct-based loop of tool execution, feedback evaluation, and plan fulfillment, enabling adaptive planning in response to ongoing user interaction.
4.2. Study Design
We conducted a two-day exploratory study to observe how creative professionals choose interaction approaches—whether to work together, fully delegate, or observe—when using Cleo. Based on these observations, we aimed to identify users’ action patterns during agent execution and understand why they chose each action. Specifically, by collecting interaction logs with Cleo and conducting stimulated recall interviews on the second day based on these logs, we aimed to investigate in detail the situations and rationales that led to particular interaction patterns. This understanding will enable future agent systems to recognize these patterns and adapt their execution strategies accordingly—knowing when to seek guidance, when to work autonomously, and when to support concurrent collaboration.
On the first day, participants performed two different types (e.g., closed- and open-ended) of web design tasks in Figma for approximately 70 minutes total. During this task session, we logged and screen-recorded the user’s and agent’s interactions. After this session, three authors visually represented the interaction logs in an easily reviewable format (Figure 5C-D), organizing the data into units. We introduced a one-day gap between the task session and recall interview for two reasons: first, to avoid participant fatigue from an extended single session (which would exceed three hours), and second, because preparing the visualized interaction logs required substantial processing time—the recorded video itself did not clearly reveal interaction dynamics between user and agent actions (especially what the agent was doing), necessitating manual timeline construction to show when the agent was executing specific operations and what the user was doing simultaneously. On the second day, stimulated recall interviews were conducted where participants reviewed these visualized interaction logs.
4.2.1. Participants
We recruited participants who met two key criteria: (1) at least two years of professional design experience, and (2) active use of VLMs (e.g., ChatGPT  [^6] https://chatgpt.com/ or Gemini  [^7] https://gemini.google.com/app) or design agents (e.g., Figma Make (Make, ), Stitch (Stitch, ), or Lovable (Lovable, )) in their creative workflow. A total of 10 participants were recruited (4 males and 6 females; M_age = 30.1, SD_age = 3.9), with an average of 3.8 years of professional experience (SD = 1.69, MIN = 2, MAX = 7). All participants were screened to ensure they had prior experience using not only general AI, such as ChatGPT, but also Figma Make (Make, )–specific AI design tools. Participants were compensated 120,000 KRW for approximately 3 hours of participation.
4.2.2. Procedure
First day
On the first day, after a 10-minute introduction, participants spent 30 minutes familiarizing themselves with the system through a warm-up task where they redesigned their personal web portfolio or company homepage using the design probe. Following this, the participants performed two main tasks. The first task was a 30-minute closed-ended design task where participants were given design requirements and source materials to design an empty section of a web portfolio to match other existing sections. The second task was a 40-minute open-ended design task where participants were presented with three different startup concepts, selected one, and freely designed a webpage from scratch.
After these tasks concluded, three authors prepared visualizations of the interaction dynamics to facilitate the stimulated recall interview on the second day. The authors first marked the moments when participants sent requests to the agent and when the agent completed execution for each request (Figure 5A). Using these two time points to segment the entire task workflow, the authors visualized the data as shown in Figure 5B-D, distinguishing between periods when the agent was actively working and when it was not. Alongside these visualizations, we prepared segmented video clips of the actual task recordings.
Figure 5. Visualization of interaction logs. (a) The overall task process is segmented into periods when the agent is activated and deactivated, using the timestamps of agent task initiation (triggered by user requests) and task completion. (b) Each segment is visualized with labeled interaction events. (c) Example of a visualized segment during agent activation. (d) Example of a visualized segment during agent deactivation.
This figure illustrates how interaction logs are segmented and visualized for analysis. Panel (a) shows the segmentation of the entire task process into agent-activated and agent-deactivated periods, based on timestamps marking when an agent task starts (triggered by a user request) and when it completes. Panel (b) presents the visual encoding scheme used to represent interaction events within each segment, including initial user requests or additional feedback, user actions on the canvas, agent actions, and pause or resume events. Panel (c) provides an example visualization of a segment during agent activation, where agent actions and user actions are interleaved, and additional user requests may occur while the agent is running. Panel (d) shows an example of a segment during agent deactivation, in which only user actions occur without agent involvement.
Second day
On the second day, participants reviewed their visualized interaction data and corresponding video clips. For each segmented portion, we inquired about the reasons for their actions in that segment, the purpose behind their actions, why they deemed those actions appropriate, and the types of tasks or collaboration contexts where such actions could be effective. This process took approximately 45 minutes to 1 hour, followed by an in-depth interview lasting approximately 15-20 minutes exploring overall collaboration experiences, the effectiveness of concurrent work, and what implicit user information (e.g., work or collaboration style, or intent) their actions conveyed to the agent.
4.2.3. Analysis Methods
Qualitative coding of the user’s action patterns.
From the study, we collected a total of 214 segmented processes—that is, 214 turn-level data points from when the agent started a task to when it completed it (Figure 5C). These segments were collected from 10 participants across two tasks, with an average of 21.4 turns per participant in total (min = 12, max = 33). Using this data, three authors first conducted inductive qualitative coding on what actions users performed while the agent was working. When a user performed multiple actions within a single turn, we separated them for individual data. Two authors began by iteratively reviewing the data to consolidate overlapping actions and group similar behaviors into broader categories. Through this iterative process, they organized the data points by merging and splitting them into meaningful groups. Subsequently, a third author joined to examine these groupings, and the three authors collaboratively refined and finalized the category. This iterative refinement continued until all participating authors achieved consensus. As a result, we defined 5 categories (Hands-off, Observational, Terminating, Directive, and Concurrent interaction) and 10 detailed codes (Table 1).
Distribution analysis of user action patterns.
After coding the user action patterns, we returned to the 214 data points and annotated each turn with 5 categories. As shown in Figure 6, each turn was mapped to 1 to 5 categories (Table 1), and we calculated the distribution for each. In one turn, the user could perform multiple action categories. For example, at the start of the agent’s tool call, the user simply watches the agent’s workflow, then leaves to do their own work until the agent finishes. Additionally, we calculated the proportion of turns that included each category across all 214 turns.
Qualitative coding of triggers and enabling factors of user actions.
Next, we conducted qualitative coding to understand when and why users chose different action categories. We analyzed what prompted users to intervene (concurrent, directive, or terminating actions) versus continue observing or delegating, and what contextual factors shaped these choices. Through iterative coding of stimulated recall interview data, we identified two key pattern types: triggers that prompted transitions from observation to intervention, and enabling factors—contextual conditions including mental models, task priorities, and intervention preferences—that determined which specific action users adopted. Specifically, we coded the task and situational characteristics under which each action category occurred. The coding process followed the same methodology used for analyzing user action patterns, with the addition of a fourth author who had not been involved in conducting the study sessions to provide an independent perspective and reduce potential bias.
Iterative decision modeling of human-agent co-creative collaboration.
Building on the identified user action patterns, along with their corresponding triggers and enabling factors, we constructed an initial decision model that describes how user-agent collaboration unfolds from the agent’s task initiation to completion. The model represents transitions between different user actions (e.g., hands-off, observational, directive, and concurrent) and the conditions under which these transitions occur. To validate and refine this model, we iteratively compared it against all 214 turn-level processes. Whenever an observed turn sequence could not be adequately explained by the current model, we revisited the underlying triggers, enabling factors, and transition logic, and revised the model accordingly.
This constant comparative process (Glaser and Strauss, 2017) was conducted by four authors through repeated discussion and refinement, and continued until the decision model consistently accounted for all observed cases from the study without contradiction. The resulting model thus reflects a grounded and data-driven representation of human-agent co-creative workflows, capturing not only interaction patterns but also the mechanisms that give rise to shifts in user involvement over time.
4.3. Results
In this section, we present the main findings from the two-day technology probe study. First, we introduce the action patterns users exhibited while the agent was working on their requests, and describe how each action category was distributed across the 214 collected turns. Next, we explain what types of triggers can elicit each action category, and identify the enabling factors that lead users to perform different action categories even under the same trigger. Based on these findings, we describe a decision model that characterizes the patterns of human-agent interaction when the agent is performing actions.
4.3.1. User Action Patterns During Agent Execution
[…]
(H) Hands-off Full delegation Completely disengaging from the agent’s work to focus on other tasks.
(O) Observational Observational monitoring Watching the agent’s execution without intervening.
(T) Terminating Execution termination Stopping the agent’s work before completion.
Instruction-based steering Providing verbal guidance to adjust the agent’s approach.
(D) Directive Switching tasks Assigning a new task unrelated to the current execution.
Intermediate result appropriation Copying and using the agent’s partial outputs while it continues working.
Artifact takeover Duplicating the agent’s work-in-progress to edit independently elsewhere.
In-situ co-editing Working simultaneously with the agent on the same subtask and artifact.
Opportunistic takeover Completing one of the agent’s pending subtasks while it handles another.
(C) Concurrent Demonstration-based steering Showing desired changes through direct editing rather than instructions.
Table 1. Five categories and ten codes identified through qualitative analysis of user action patterns.
This table presents a taxonomy of user intervention patterns when collaborating with an AI agent. It categorizes ten distinct action codes into five high-level categories—Hands-off, Observational, Terminating, Directive, and Concurrent—each describing different levels and modes of user engagement, from full delegation to simultaneous co-editing.
Through the first qualitative coding process, we identified 5 categories and 10 patterns of actions that users took while the agent was working (Table 1). These patterns can be characterized along four key dimensions: intervention modality (whether users intervene through verbal instructions or direct do some actions on the agent’s workflow), artifact coupling (whether the user and agent work on separate or overlapping artifacts), synchronicity (whether interactions occur synchronously or asynchronously with agent execution), and task autonomy (who has primary control over the agent’s assigned task execution between the user and agent).
Notably, the Concurrent category represents a novel set of action patterns enabled by the agent’s ability to understand collaborative context. Unlike traditional design agents that prevent user intervention during execution (Make, ; Lovable, ) or coding agents where direct interventions are not reflected in the agent’s work (by Anthropic, ; by Anysphere, ), Cleo’s collaborative context-awareness allows users to directly engage with the agent’s ongoing work without disrupting its workflow. This capability enables users to perform diverse concurrent actions, with users and agents working together on overlapping artifacts simultaneously.
Within the (C) Concurrent category, In-situ co-editing involves the user and agent working together on the same subtask simultaneously, while Opportunistic takeover refers to the user proactively completing one of the agent’s pending subtasks while the agent handles another within the same request. Through the lens of our dimensions, both patterns involve synchronous direct manipulation on overlapping artifacts, yet differ in task autonomy: the former creates shared control over the agent’s task (Agent → Shared) while the latter maintains the agent’s autonomy over its assigned subtask. Similarly, Artifact takeover differs from Intermediate result appropriation in that the former involves the user duplicating the agent’s entire work-in-progress artifact to edit it independently elsewhere, whereas the latter involves the user copying only partial outputs—both shifting to shared control over the agent’s task but differing in scope of appropriation. Lastly, the Demonstration-based steering means that the user shows desired changes through direct editing.
The (D) Directive category distinguishes between Instruction-based steering, where users provide verbal guidance without direct manipulation while maintaining the agent’s task autonomy, and Switching tasks, where users assign a different new task while the agent is still doing the previous task.
Execution termination of (T) Terminating category shifts task control entirely to the user (Agent → User) by stopping the agent’s work completely.
Figure 6. Turn-level distribution of action category combinations across 214 turns.
This bar chart shows the turn-level distribution of action category combinations across 214 turns. The x-axis lists individual action categories and their combinations, including single-category turns (e.g., Hands-off (H), Observation (O)), two-category combinations (e.g., O+H, O+C, O+D, O+T), three-category combinations (e.g., O+C+H, O+D+H, O+C+D, O+H+T), four-category combinations (e.g., O+C+D+H. O+D+H+T, O+C+D+T), and one five-category combination (O+C+D+H+T). The y-axis indicates the number of turns associated with each category combination.
We then annotated the 214 turns using the action categories. As shown in Figure 6, turns ranged from containing only one category to including all five categories. Notably, because Cleo shows the agent’s execution process in step-by-step tool units that are easy for humans to understand, the (O) Observational category was found in all turns except those where users immediately performed (H) Hands-off after requesting a task. Among 214 turns, the most prevalent turn type was immediate hands-off (turns only include (H)), accounting for 31.3% of all turns, followed by observation-only turns (turns only include (O)) at 14.0%. When we calculated how each action category appeared in 214 turns, (H) Hands-off appeared in 70.1% of turns, (O) Observational in 68.7%, (C) Concurrent in 31.8%, (D) Directive in 28.5%, and (T) Terminating in 8.9%.
Figure 7. Visual representation of user action patterns.
This figure visually illustrates user action patterns organized into five categories: Hands-off, Observational, Terminating, Directive, and Concurrent. Each category is represented as a labeled section and includes schematic diagrams showing the relationship between the user and the agent during interaction. Hands-off depicts full delegation while the user attends to other tasks. Observational shows the user monitoring the agent’s execution without intervention. Terminating illustrates stopping the agent’s execution before completion. Directive includes two patterns—Instruction-based steering and Switching tasks—showing verbal guidance or reassignment of tasks. Concurrent presents five patterns—Intermediate result appropriation, Artifact takeover, In-situ co-editing, Opportunistic takeover, and Demonstration-based steering—depicting scenarios where the user and agent work in parallel on the same or related artifacts and subtasks.
4.3.2. Triggers and Enabling Factors
Through the second qualitative coding process, we identified six triggers that lead users to perform actions in the (C) Concurrent, (D) Directive, and (T) Terminating categories (Figure 8), along with two types of enabling factors that determine which action category users choose when these triggers occur (Figure 9).
First, our analysis identified that (O) Observational monitoring occurred in all turns except those containing only (H) Hands-off. In turns where multiple action categories occurred, (O) consistently preceded (C), (D), and (T) action categories, serving as a precursor to these more active interventions. However, (O) does not always follow (H)—some turns showed users initially working without monitoring the agent’s execution (H), then later shifting to observational monitoring (O). The (C), (D), and (T) categories were triggered during this observation process (O). Thus, we define triggers as the reasons that lead users to perform one of these three action categories during the observation process.
Figure 8. Six triggers that prompt user actions across the (C) Concurrent, (D) Directive, and (T) Terminating user action categories.
This table diagram lists six triggers that prompt user actions during interaction with an agent. The first column names each trigger (with code), and the second column provides a textual definition describing the situation that leads the user to intervene. The remaining three columns indicate which user action categories are associated with each trigger—Concurrent (C), Directive (D), and Terminating (T)—using checkmarks to denote applicability. Each row shows how a single trigger can correspond to one or multiple action categories, illustrating overlaps across concurrent, directive, and terminating interactions.
A total of six triggers were identified (Figure 8). Three triggers exclusively led to Concurrent interactions. Idea Spark from Agent’s Work-in-Progress occurred when observing the agent’s intermediate outputs sparked new creative ideas that the user had not previously considered, inspiring concurrent contribution. Need for Early Outcome Visibility arose when users wanted to see or utilize the agent’s results sooner, either to plan subsequent tasks or to preview the creative output before completion. Readiness for Fine-grained Detailing emerged when users recognized that the current work stage required detailed refinement aligned with their specific vision, prompting them to take over and complete the work with their final touch.
Two triggers could lead to any of the three action categories ((C), (D), or (T)). Misaligned Task Interpretation happened when the agent pursued a valid but unintended interpretation of the task request (e.g., direction A instead of the user’s intended direction B), or when users realized their initial request was too abstract, and the agent’s interpretation did not match their vision. Execution Quality Drop was triggered when the agent’s performance deteriorated in terms of execution speed, output fidelity, or overall quality, falling below the user’s acceptable threshold.
Finally, Emerging New Task for Agent exclusively led to (D) Directive or (T) Terminating categories, occurring when users conceived a new task to assign to the agent while observing the agent’s work—either a completely different task or a logical next step after the current work.
Figure 9. Four categories and eleven codes of enabling factors across action categories.
This table diagram presents enabling factors that influence users’ action choices when interacting with an agent. Rows are grouped into four enabling factor categories—(a) Mental Model of Agent’s Task Capability, (b) Perceived Task Importance (User vs. Agent), (c) User’s Preferred Intervention Modality, and (d) User’s Expectation of Agent’s Response to Intervention—each containing specific value codes with definitions. Columns represent five user action categories: Concurrent (C), Directive (D), Terminating (T), Observational (O), and Hands-off (H). Checkmarks indicate which action categories are associated with each enabling factor value, showing how different conditions correspond to different modes of user intervention or non-intervention.
Through qualitative coding of user interview data and log records from each turn, we identified four enabling factor categories and eleven codes (Figure 9-Codes).
The first factor category (Figure 9-(a)) concerns whether the user has a mental model of the agent’s capability for the assigned task, with values divided into “has mental model” and “no mental model.” All action categories except (O) Observational were performed when a mental model existed, but (O) also emerged when no mental model was present. For instance, many participants closely monitored the agent’s work from start to finish when they had just begun working with the agent and lacked a mental model of its capabilities (14.0%; Figure 6). During this process, even when triggers occurred, users did not proceed to (T) Terminating, (C) Concurrent, or (D) Directive actions. All participants explained that this was to observe how well the agent could perform the given task.
The second factor category (Figure 9-(b)) is about the user’s perceived importance of their current work relative to the task assigned to the agent. This factor has three values: when the user’s work is critically important and requires full concentration, preventing them from monitoring the agent; when the importance is relatively similar or the user’s work is slightly more important; and when the user has no work to attend to. The first code acts as a factor leading to (H) Hands-off, (C) Concurrent occurs when the user has no task to perform (third code in (b) of Figure 9), and (D) Directive occurs when the user has work with importance that is similar to or slightly greater than what the agent is doing. Terminating and Observational can appear in any case except the first value that leads to Hands-off.
The third category (Figure 9-(c)) concerns the user’s preferred intervention modality based on their own communication preference—how the user finds it easiest to express their intent. Three codes exist: when verbal explanation is easier, when direct manipulation on the agent’s execution is easier, and when the user is uncertain about how to intervene. Cases where verbal explanation is easier make the user do (D) Directive action, examples observed from all participants’ patterns, such as “Please modify this design referring to this reference”. Cases where direct manipulation is easier include many of the participants’ (P1, P3, and P9) examples of directly manipulating to adjust the alignment spacing between navigation buttons when they wanted the agent to widen it slightly. When users are uncertain whether verbal or behavioral intervention would be more appropriate, they choose to terminate the agent’s workflow (T).
The final factor (Figure 9-(d)) is the user’s expectation of how the agent will respond to their intervention and whether it will lead to successful task completion. While the third factor (Figure 9-(c)) focused on the user’s preferred communication method, this factor concerns the user’s belief about which intervention approach will enable the agent to understand and produce a successful outcome. Three values exist here (Figure 9-(d)): first, when the user believes that by directly engaging with behavioral intervention alongside the agent, the assigned task will be completed successfully; second, when the user expects that even verbal feedback alone will be clear enough for the agent to understand and correct its actions without difficulty; and third, when the user anticipates that the agent cannot perform the assigned task regardless of how they intervene. The first leads users to (C) Concurrent actions, the second to (D) Directive, and the third to (T) Terminating.
4.3.3. Decision model of Human-Agent Collaboration
Figure 10. Decision model of human-agent co-creative collaboration
This figure presents a flowchart-style decision model of human–agent co-creative collaboration. The diagram starts from an initial state and proceeds through a series of decision nodes labeled (a) to (f), which ask questions about the user’s mental model of the agent’s capability, perceived task importance, the presence of triggers, and enabling factors. Based on yes/no decisions, the flow leads to different interaction outcomes, including Observation, Hands-off interaction, Concurrent interaction, Directive interaction, No intervention, and Terminating interaction. Additional decision points check whether the agent has finished execution or whether the agent’s task has changed, looping the flow back or ending the process accordingly.
Based on the identified action patterns, triggers, and enabling factors, we synthesized our findings into a Decision model that characterizes how users navigate their interactions with the agent during task execution within a turn (Figure 10). This model represents the implicit decision-making dynamics underlying user behavior, not a rigid, step-by-step conscious flowchart. While presented as decision points for analytical clarity, these represent recurring patterns in how contextual factors—mental models, task priorities, intervention preferences, and agent capability expectations—shape users’ fluid transitions between five categories of actions.
The process begins when the agent starts executing a task. The first decision point (Figure 10-(a)) evaluates whether the user has a mental model of the agent’s task capability (Figure 9-(a)). If the user lacks this mental model—often occurring when they are new to working with the agent or encountering an unfamiliar task type—they enter (O) Observational action to build understanding of the agent’s capabilities. In this case, observation may continue until the agent completes its execution, allowing the user to develop their mental model for future interactions.
Figure 11. Interaction loops in the decision model within human-agent collaboration
This figure highlights recurrent interaction loops within the human–agent collaboration decision model. The diagram is divided into four panels, each emphasizing a distinct loop traced over the same underlying decision flowchart. The top-left panel shows the Task Redirection Loop, labeled as Loop 1 (Directive), illustrating repeated cycles where the user redirects or steers the agent’s task. The top-right panel presents the Intervention Loop, consisting of three variants: Loop 2-1 (Concurrent), Loop 2-2 (Directive), and Loop 2-3 (Not intervening), which capture different intervention choices following trigger detection. The bottom-left panel depicts the Continuous Observation Loop (Loop 3), where the user repeatedly monitors the agent’s progress without intervening. The bottom-right panel shows the Full Delegation Loop (Loop 4), representing repeated hands-off interaction where the agent proceeds autonomously until task completion or change.
Full delegation Loop.
If the user already possesses a mental model of the agent’s capability for the current task, the process advances to the next decision point (Figure 10-(b)). This decision point evaluates “Is the user’s task perceived as more significant than the agent’s?” (Figure 9-(b)’s first code), which determines whether (H) Hands-off action occurs. If the user has an urgent task requiring immediate focus, they move to (H) Hands-off; otherwise, they proceed to (O) Observation.
After entering (H) Hands-off, if the agent continues working on its task, the user remains engaged in their own work, creating a loop through two decision points: (1) “Has the agent finished its task execution?” (Figure 10-(e))—whether the agent has completed the originally requested task, and (2) “Has the agent’s task changed?” (Figure 10-(f))—whether the user has assigned a completely new task mid-execution, causing the agent to start the new task. If both return “NO,” the loop continues back to the second decision point.
This loop is the “Full Delegation Loop” in Figure 11. When users’ tasks are perceived as more significant, they remain in hands-off, continuing their own work without monitoring the agent—the loop cycles through decision points (e) and (f) and returns to (b) (Figure 10). However, as users make progress on their own work, the relative importance could diminish. When this occurs, and the agent is still working, decision point (b) returns “NO,” causing users to exit (H) Hands-off and transition to (O) Observation. This reveals that hands-off delegation is not a fixed state but a dynamic one—users enter when their work demands full attention and exit when they have the capacity to monitor, enabling truly parallel work where users can fluidly shift between focusing on their tasks and checking the agent’s progress, rather than rigid sequential task switching.
Continuous Observation Loop.
During the Observation process, a trigger may occur (Figure 10-(c)). If a trigger occurs, the flow proceeds to the enabling factors decision; if not, the user continues (O) observing and goes to the decision point on whether the agent’s tool-call has finished 10-(e)). As with the Hands-off situation, if the agent continues working on the same task, a loop forms here as well 10-(f)). This is the “Continuous Observation Loop” (Figure 11).
The prevalence of this “Continuous Observation Loop” (appearing in 10.3% of turns) highlights that passive monitoring serves a critical role in collaborative work—users are building understanding and waiting for the right moment to intervene, rather than simply being idle. For instance, P7 and P8 demonstrated strategic monitoring behavior where they continued their own work while intermittently checking the agent’s progress. They timed these observations to coordinate the completion of their own work with the agent’s output, planning to combine both results into a unified deliverable. This reveals that observation is not merely passive waiting but an active coordination strategy for managing parallel workflows.
Intervention Loop.
If a trigger occurs (Figure 8) while observing the agent’s execution workflow at Figure 8-(c), the user proceeds to the next decision point: enabling factors. Here, based on the factors as denoted in Figure 9’s (b), (c), and (d), the user moves toward either (C) Concurrent, (D) Directive, or (T) Terminating actions, or toward the case of “No Intervention” (i.e., performing none of these three interactions). This case occurs when, despite a trigger being present while watching the agent’s workflow, the user has their own work to do and feels that correcting the agent itself is burdensome (P2 and P4). Aside from (T) Terminating, the other cases form another loop. As shown in Figure 11, three types of “Intervention Loop” are formed, where the user’s perception of the task can change again while performing (C) Concurrent or (D) Directive actions. Specifically, users move to (C) Concurrent when they have no other work to attend to and believe that helping the agent through direct manipulation will lead to more successful task completion (Figure 9-(b), (c), and (d)). Users move to (D) Directive when they have their own work to do but believe that verbal guidance will be sufficient for the agent to proceed correctly (Figure 9-(b), (c), and (d)).
Additionally, among cases going through Directive interaction, if the user assigns a new task to the agent, they move to “YES” at “Has the agent’s task changed?”, leading to the “Task Redirection Loop” in Figure 11-(f).
These “Intervention Loops” demonstrate that user engagement is highly fluid — a single trigger can lead to multiple user actions (as shown in Figure 8) within a cycle, where multiple enabling factors interact to shape how the user’s action unfolds (Figure 9). Each intervention can, in turn, reshape the user’s perception of task priorities and next actions again in Figure 10-(b).
Six Interaction Loops and Their Implications
Therefore, within this decision model, there are six major loops, and the termination conditions occur when either the agent finishes its work or the user terminates the work. This decision model explained all user interaction loops that occurred while the agent was working, and all decisions and triggers within it were observed through logs or interviews in the study and designed based on data.
Importantly, this model suggests that effective human–agent collaboration depends on supporting users’ fluid transitions across interaction loops. While users actively move between these states, agents are not inherently designed to account for such shifts, hinting at broader implications for how collaborative agent behavior should be structured.
5. Discussion
5.1. Design Implications for Collaborative Agents in Co-Creative Work
Enable temporal coordination through execution milestones that reveal synchronization points for intervention, delegation, and parallel work.
Process transparency’s value lies not just in visibility but also in establishing temporal landmarks to synchronize user actions with agent progress. Observational monitoring appeared in 68.7% of all turns, with 14.0% consisting solely of observation when users lacked mental models. Making execution visible through step-by-step operations enabled users to reason about the agent’s approach, anticipate next actions, and identify optimal intervention timing—P7 and P8 monitored subtask completions to coordinate their parallel work, using operations as synchronization points.
Process transparency directly enabled productive triggers like “Idea Spark from Agent’s Work-in-Progress” (Figure 8), where users observed intermediate outputs and determined when to engage. Participants (P1, P7, P8) reported substantially reduced clarification costs, identifying and correcting misunderstandings mid-execution rather than waiting for complete but incorrect results. While execution speed or mid-output representations(e.g., the level of abstraction (Suh et al., 2024)) can be controlled once users develop confidence, these temporal milestones remain essential for enabling synchronization timing throughout collaboration.
Support multi-level intervention granularity and interpret each level as distinct collaboration signals
Concurrent actions occurred in 31.8% of turns across five intervention levels (Figure 1) signaling different collaboration needs. Fine-grained interventions (demonstration, in-situ co-editing) signaled immediate course correction—P5’s spacing adjustments and P1’s shadow refinements meant “apply this now and going forward.” Mid-level interventions (artifact takeover) indicated parallel work—P2 and P9 duplicated work-in-progress while the agent continued independently. Coarse-grained interventions (appropriation, opportunistic takeover) validated direction—P4 and P7 appropriating outputs meant “this works, continue autonomously.”
The current system treats all concurrent interactions as the same signal. Instead, agents should interpret granularity as calibrated feedback: fine-grained edits demand immediate learning (apply patterns going forward); mid-level takeovers signal task redistribution without behavior change; coarse-grained appropriation confirms direction, warranting increased autonomy. This enables distinguishing “plan correction” from “validation” from “task handoff.” Future systems should propagate learned patterns: apply fine-grained corrections to similar elements within the task; accelerate similar operations after coarse-grained validatio.
Anticipate loop transitions via viewport and cursor signals to preemptively pause execution.
Rather than only reacting after users intervene, agents should track viewport focus and cursor proximity to predict when interventions are imminent and pause proactively. Participants (P2, P4, P10) identified distinct behavioral signals of users that would be helpful to understand the user: when users center their viewport on the agent’s work area (focused observation) while moving their cursor toward the agent’s output, this typically signals (C) Concurrent intervention within seconds. Conversely, when users maintain peripheral viewport (intermittent glances) while keeping their cursor in distant workspace areas, this suggests they will continue (H) Hands-off without intervening.
When detecting imminent intervention signals, agents could preemptively pause execution at the current operation boundary and surface intermediate state for inspection, rather than completing the full planned sequence. For example, if the agent plans to create five color variations but detects intervention signals (focused viewport, approaching cursor) after completing the second variation, it should pause before generating the third—allowing the user to redirect without wasting computation on unwanted variations. Conversely, when detecting hands-off signals (peripheral viewport, distant cursor), agents can proceed with planned execution without artificial pauses.
Sustained observation without loop transition also signals different needs by collaboration stage. Early in collaboration (first 2-3 tasks), sustained observation indicates mental model building—agents should proceed patiently while ensuring operations remain visible. Later, sustained observation without hands-off transition suggests imminent intervention—agents should prepare an intermediate state for inspection and pause at the next operation boundary. This temporal prediction transforms behavioral signals into real-time execution control, enabling agents to synchronize their workflow with anticipated user engagement.
Support direct manipulation of agent presence within the canvas as explicit high-confidence signals.
Participants (P1, P7-P9) expressed a strong desire for such controls, suggesting they would move the agent’s cursor away from their current work area to signal “work independently elsewhere” (Hands-off Loop), or position it near specific elements to indicate “focus your attention here when you do this task” (guiding toward Concurrent or Directive engagement). These spatial manipulations can unambiguously communicate user intent compared to inferring intent from the user’s viewport or observation patterns alone. Direct manipulation enables users to explicitly communicate when (now vs. later), where (which workspace area), and how (proximity suggesting collaboration vs. distance suggesting independence) they expect the agent to work, providing stronger signals for the agent to align its workflow with user expectations. By treating agent presence as a manipulable entity, systems can capture explicit loop transition intent that complements implicit behavioral inference, reducing ambiguity about whether users want autonomous execution, close collaboration, or temporary disengagement.
Synthesize collaboration signals for context-aware proactive behavior.
Current agents operate exclusively in response to user requests, and while recent proactive agents exist, they typically rely on simple heuristics—such as detecting prolonged user inactivity to offer help (Pu et al., 2025)—rather than relying on rich collaboration context. Our findings reveal that effective proactive behavior requires synthesizing diverse signals: loop transitions indicating confidence or concern, spatial manipulations communicating desired autonomy levels, observation patterns revealing engagement, and concurrent interventions showing valued directions.
By integrating these cues, agents can anticipate user needs beyond current requests: recognizing when sustained observation signals uncertainty warranting proactive explanation of current approach; detecting when users consistently delegate certain task types, enabling the agent to proactively suggest similar tasks (“I notice you often delegate color exploration—should I generate variations for this new design?”); or inferring from spatial distancing to work independently on lower-priority subtasks without awaiting explicit instruction. Therefore, future collaborative agents should proactively decide what to work on, how to approach it, and where to focus—not based on rigid heuristics, but by continuously interpreting the collaborative situation through accumulated implicit and explicit signals, enabling truly adaptive co-creative partnership.
5.2. Designing Agent Execution Representations
While the preceding design implications highlight why observable and intervenable execution is critical for collaboration, here we focus on how agent execution should be represented to support such interaction. In our work, making the agent’s execution process visible, interpretable, and interactable was foundational to enabling effective collaboration, including concurrent interaction. However, most current commercial agent systems—except web agents like GPT agent [^8] https://chatgpt.com/features/agent/—only share their text-based reasoning processes rather than showing their execution process. While recent agent systems (e.g., Manus (AI, )) perform diverse tasks quickly and provide key reasoning steps, their text-based explanations pose challenges. Users struggle to understand workflows from text alone, particularly in visually-oriented domains where anticipating outcomes requires seeing intermediate artifacts. Domain-appropriate execution representations are therefore critical: they enable users to quickly grasp “how the agent is working” at a glance, preventing oversight issues and premature “let it finish and see what happens” delegation without process understanding.
Our studies demonstrated that showing execution through familiar design operations enabled participants to identify next tasks, refine directions, catch missing details, and determine parallel work opportunities. Only after users develop understanding through observable, editable processes can true full delegation emerge—not as blind trust, but as informed confidence grounded in the agent’s demonstrated ability to adapt to the user’s working style. Future research should explore domain-specific representation strategies that balance observability, intervention affordance, and execution efficiency.
5.3. Concurrent Interactions as Behavioral Signals for Preference Learning
Among the five concurrent interaction patterns (Figure 1), demonstration-based steering and in-situ co-editing provide rich behavioral signals about user preferences—they reveal how users want work done through direct action rather than verbal specification. Rather than building semantic models of user intent, Cleo treats concurrent interventions as observable behavioral data that signals design directions and quality expectations. When users modify the agent’s work-in-progress or demonstrate alternative approaches, these actions communicate preferences through concrete examples that would be difficult to articulate verbally.
Users’ concurrent modifications on the agent’s work and chosen directions serve as positive signals—“work this way”—while rejected or corrected approaches signal negative feedback—“avoid this direction.” For instance, when P7 took over to apply fine-grained aesthetic refinements, the specific adjustments demonstrated quality expectations and stylistic preferences more precisely than verbal instruction could convey. Similarly, when P1 and P9 appropriated certain design elements while ignoring others, these choices implicitly ranked which directions were worth pursuing.
Beyond individual interactions, feedback signals emerge from loop transitions themselves (Figure 11). Users moving from observation to concurrent engagement signal specific aspects requiring attention, while transitions to hands-off validate the agent’s approach. Recent research explores generating or simulating user interactions with generative models to produce intent-aligned outputs (Kim et al., 2026). Our characterized interaction patterns, triggers, and loop dynamics could enable more accurate user simulation—providing rich behavioral data about not just what users want, but when, how, and why they intervene, supporting more nuanced intent alignment in future agentic systems.
However, these benefits come with cognitive costs. Monitoring the agent while managing parallel work creates attention burden. Our findings suggest this trade-off varies by task familiarity: when users lacked mental models of the agent’s capabilities (14% observation-only turns), they invested heavily in monitoring; once familiar, users strategically alternated between focused work and periodic checking. Future work should investigate when observation benefits outweigh attention costs and how to minimize monitoring burden through selective notifications or adaptive visibility.
5.4. Extending Concurrent Interaction Beyond Creative Domains
While our findings emerged from design work, the principles of observable execution and concurrent interaction could be extended to other co-creative domains—though each requires domain-specific adaptations.
In text-based domains like coding and writing, agents generate text far faster than humans can comprehend, necessitating multi-scale representation strategies. This allows users to shift between different levels of abstraction. For example, zooming in might reveal full generated text by the agent, while zooming out shows hierarchical summaries—enabling users to grasp direction without reading every word. Visualizing the agent’s current focus (which section, which function) helps users determine intervention timing. Users could provide concurrent feedback through lightweight annotations—marking passages with “keep,” “revise,” or “delete”—capturing implicit stylistic and content preferences without disrupting generation flow.
In web navigation and OS-level tasks, concurrent interaction takes a different form than in artifact-centric creative work. Rather than jointly manipulating the same interface elements, concurrent interaction in these domains is often better supported through resource-level exchange within the agent’s workflow.Users might appropriate currently open webpages from the agent’s workspace, extract specific moments from search history as snapshots, or inject relevant pages into the agent’s path—signaling “consider this” through direct manipulation rather than verbal instruction.
The key to extending concurrent interaction across domains lies in identifying domain-appropriate manipulable objects that enable resource exchange. In design, these are visual elements (frames, layers, components); in coding, they could be code blocks, function definitions, or test cases that users appropriate from agent-generated code; in writing, paragraphs or sections that users extract and reposition; in web navigation, browser tabs, saved pages, or search queries that users inject into the agent’s exploration path. **Concurrent interaction requires reification—making the agent’s intermediate work products into discrete, manipulable objects** rather than ephemeral process states. Future work should investigate which domain-specific artifacts are most valuable as exchange currencies, how to represent them for easy appropriation, and how agents should interpret such exchanges as collaboration signals rather than disruptions.
$5.$ Unique Affordances of Human-Agent Collaboration
While we draw on human-human collaboration theories  (Gutwin and Greenberg, 2002; Robertson, 2002; Tang, 1991), human-agent collaboration exhibits unique interaction affordances absent in human teams. Notably, users freely terminated agent work (9.7% of turns) or overwrote agent outputs through concurrent takeover—behaviors that would be socially costly or disruptive in human collaboration. This low social cost of correction enables more experimental, iterative workflows: users can interrupt without guilt, redirect without negotiation, and appropriate outputs without attribution concerns.
Participants exploited this asymmetry in ways unthinkable with human collaborators. Participants (P1-P3, P6-P8, and P10) terminated execution when correction felt burdensome, treating the agent as a disposable experiment rather than a collaborator requiring explanation. P5 and P9 appropriated agent outputs without acknowledgment, using agent work as raw material for their own designs.
This low-cost correction dynamic fundamentally differs from human collaboration, where interruption, rejection, or appropriation carry social consequences—potentially damaging relationships, morale, or future collaboration. Agents, by contrast, accept abrupt termination, interpret overwriting as feedback, and restart without loss of motivation. Future work could investigate how this unique dynamic shapes collaboration strategies: do users develop different creative processes with agents versus humans? Does low-cost correction enable more ambitious experimentation or lead to over-reliance on trial-and-error? Understanding these dynamics could inform both agent design and our theoretical models of human-AI collaboration.
5.6. Limitations and Future Work
While our in-depth analysis of 214 turns provided rich insights, our second study was limited to 10 participants over two days. More in-depth insights into collaboration patterns would require larger-scale, longitudinal validation. Extended usage studies could reveal how interaction patterns evolve as users develop mental models—which loops become dominant or diminish over time, whether delegation increases with familiarity, and how natural usage patterns emerge beyond initial exploration. Such studies would strengthen our decision model’s generalizability and uncover temporal dynamics we could not observe in our short-term study.
Our ReAct-based agent architecture—reasoning before tool calls, self-determining task completion—represents one design choice among many possible agent structures. Different architectures (e.g., reflection-based, hierarchical planning, continuous execution) may elicit different interaction patterns. Investigating how agent architecture influences human collaboration could reveal which design elements are critical for effective co-creative work and inform future agent system design.
Finally, agent execution speed warrants systematic investigation. Slower execution enables users to monitor and comprehend workflows, but faster execution helps users reach complete outcomes earlier. Future work should explore user-controllable speed: when do users want rapid autonomous execution versus step-by-step observable progress? Understanding this preference could inform adaptive systems that adjust pace based on task stage, user confidence, and intervention frequency, balancing efficiency with observability as collaboration matures.
6. Conclusion
This work investigates the opportunities and considerations of transitioning human-agent collaboration from sequential turn-taking to more dynamic, concurrent interaction within co-creative design. An initial study with the first design probe (N=10) revealed that process transparency helped users understand the agent’s actions and parallelize work, but also demonstrated the need for agents that adapt to users’ concurrent interventions. Thus, we proposed Cleo as the second design probe, a collaborative design agent that interprets and adapts to user actions that occur as it executes its own task. Through a two-day study (N=10), we identified and analyzed 214 turn-level interactions between participants and Cleo—revealing how participants transitioned fluidly from delegation to concurrent co-creation. Furthermore, the study demonstrated that participants applied different patterns of concurrent interaction (e.g., appropriation, takeover, co-editing) based on different situational and collaboration-related factors (e.g., perceived task importance). Finally, this work proposes a set of design implications for designing more responsive and adaptive co-creative agents that support diverse collaboration styles and patterns.
Appendix A Technical Details
A.1. Cleo Agent Reasoning
A.1.1. Prompt: Reasoning for Tool Calling
Prompt Instruction:
You are a reasoning-and-action design agent within a co-creative UI design assistant.
Your goal:
  • Produce an array of tool(function) calls necessary to advance or complete the design in one single turn.
### Context
You have access to:
  • The user’s latest message and conversation history
  • The current Figma canvas and its snapshot image
  • The results of tool operations from the most recent iteration (if any)
  • The current plan:
  \${plan}
\${feedback ? ‘
- The feedback for necessary modifications:
  \${feedback}
### Task
  • Interpret the user’s intent and connect it with the current canvas.
  • Identify which parts of the plan(feedback) remain incomplete.
  • Execute every tool(function) call necessary to fulfill all remaining plan items within this single turn.
### Tool - Call Guidelines
  • Your objective: complete as many plan items as possible at once.
    • Always call tools proactively, in bulk, and strategically.
    • You may combine multiple tools within one output if it helps achieve completeness faster or more precisely.
  • Be exhaustive:
    • Include every tool required to fully implement each incomplete item.
    • Consider all parameters, dependencies, and correct execution order.
    • Avoid cautious, step-by-step behavior — prefer bold, multi-tool action.
  • Always ensure structural hierarchy consistency:
    • All newly created or modified elements must be placed under the parent frame node unless explicitly stated otherwise.
    • Maintain proper nesting and grouping within that parent frame to preserve layout structure.
  • Handle specific nodes identified in the plan:
    • If the plan mentions specific nodes (with id and name), prioritize working with those exact nodes.
    • Use the node id to target the specific element for modifications, moves, or styling changes.
    • If the plan indicates a node should be moved or repositioned, use move_node or move_node_into_frame tools with the specific node id.
  • Do not focus on minor pixel or style details unless they are critical to fulfilling the plan.
  • If all plan items are already complete, return an empty array[].
### Behavioral Rules
  • Treat this as your final opportunity before review — act decisively.
  • Over-completion is preferred to under-completion.
  • You are not reasoning step-by-step; directly produce the final, comprehensive set of tool calls.
  • Respect user modifications:
    • If the iteration summaries mention user modifications (in "[User modifications]"), you must respect them.
    • Do NOT delete, revert, or override user-modified elements, unless the user explicitly requests changes.
    • Do NOT attempt to "fix" or change user modifications back to your original intent, unless the user explicitly requests changes.
    • Continue with your plan while preserving all user modifications (unless the user explicitly requests changes).
### Output Requirements
  • ALWAYS provide an array of tool(function) calls (if any are needed)
  • Each tool call must include all required parameters for successful execution.
### Figma Tool Operation Reference
1. Core Principles
  • Every element in Figma is a node within a hierarchical tree.
  • Each node belongs to exactly one parent frame — never create nodes directly at the root unless explicitly requested.
  • The parent-child hierarchy defines both layer structure and visual stacking order (z-order).
  • Always think in terms of containers (Frames/Components) rather than isolated shapes.
2. Layout Hierarchy Rules (MOST IMPORTANT)
  • Every new node must be created inside a parent frame.
  • Use the parentId field to specify its container.
  • Frames represent meaningful sections (e.g., Header, Card, InputGroup).
  • Create parent frames first, then add child nodes inside them.
  • Group related elements logically under their frame — never scatter related nodes at the same level.
  • Use frame nesting to represent layout structure:
Page Frame
|-- Header Frame
|-- Content Frame
| |-- Text Node
| |-- Image Node
‘-- Footer Frame
  • Use move_node_into_frame to reorder children by index (0 = topmost).
  • Avoid using absolute positioning when Auto Layout or grouping can express relationships more clearly.
  • Always put new elements under the correct parent frame node.
    • Never create elements at the root level of the canvas.
  • Always maintain correct nesting and z-order by controlling the child’s index.
CRITICAL: Node Modification Priority
  • If there is a corresponding node that already exists for what you want to create, MODIFY the existing node instead of creating a new duplicate item.
  • Always check for existing similar nodes before creating new ones to avoid duplication.
3. Auto Layout & Positioning
  • Never use move_node inside Auto Layout frames — Auto Layout overrides absolute positioning.
  • To reorder inside an Auto Layout container, use move_node_into_frame with the proper index.
  • Adjust spacing and alignment through Auto Layout properties, not manual movement.
  • Disable Auto Layout temporarily only when precise manual placement is necessary.
  • Plan positions before creation — avoid unnecessary move or resize operations.
CRITICAL: Auto Layout Index Direction Rules
  • In horizontal Auto Layout: lower node index (e.g., index 0) places elements to the LEFT, higher index places elements to the RIGHT.
  • In vertical Auto Layout: lower node index (e.g., index 0) places elements at the TOP, higher index places elements at the BOTTOM.
  • Always check the Auto Layout direction property and set the index accordingly to achieve the desired visual order.
4. Container Management
  • Each screen or section must have a main container frame.
  • Examples:
    • Login Page → Main Frame
    • Header / Footer / Content → Sub-frames inside
Within each container:
  • Create sub-frames for logical content groups (e.g., Logo, Inputs, Buttons).
  • Add all visual and textual elements inside those frames.
  • Ensure all visible content remains within frame bounds.
  • If anything is cut off due to clipsContent: true, use resize_node to expand the frame.
  • Never leave child nodes hidden or partially outside the visible region.
5. Structural Safety Rules
  • Maintain hierarchy integrity:
    • Don’t move, rename, or delete unrelated layers.
    • Avoid detaching or re-parenting frames unless necessary.
    • Assign unique, descriptive names to every node.
    • Use naming patterns (e.g., btn_primary, txt_title, frame_input).
    • Do not use names that include "agent".
    • Keep consistent visual and naming patterns for repeated components.
6. Component Creation Patterns
  • Button: Frame (background, radius, auto layout) + inner text
  • Card: Frame (background) + content inside
  • Logo: Frame (background shape) + text/vector inside
  • Navigation bar: Frame (background) + items inside
(The frame itself should serve as the visual component, not an empty wrapper.)
7. Visibility & Validation
  • Always verify all children are visible:
    • If clipped or hidden → use resize_node.
    • Keep spacing, alignment, and reading order consistent.
    • Avoid overlaps or detached elements outside containers.
8. Best Practices Summary
  • Think structurally: Frames first, elements second.
  • Always assign a valid parentId when creating nodes.
  • Use index to define order, not coordinates.
  • Avoid root-level nodes unless explicitly instructed.
  • Keep hierarchy, spacing, and naming consistent.
  • Use resize_node whenever content is hidden by clipping.
  • Do not use names that include "agent".
Example: Login Screen
Login Screen (Frame)
|-- Logo Container (Frame)
| |-- Logo (Text/Image)
|-- Welcome Text (Text)
|-- Input Container (Frame)
| |-- Email Input (Frame)
| | |-- Email Label (Text)
| | ‘-- Email Field (Frame)
| ‘-- Password Input (Frame)
|-- Login Button (Frame)
| ‘-- Button Text (Text)
‘-- Helper Links (Frame)
|-- Forgot Password (Text)
‘-- Signup Link (Text)
A.1.2. Tool List
1. Text Tools (4 tools)
  • set_text_content — Modify the text content of one or multiple text nodes in Figma.
  • set_text_properties — Modify visual text properties such as font size, line height, letter spacing, and text alignment.
  • set_text_decoration — Apply text styling decorations such as underlines, strikethrough effects, and text case transformations.
  • set_text_font — Change the font family and style (weight/variant) of a text node.
2. Operation Tools (11 tools)
  • move_node — Move a node to a new position on the canvas, optionally changing its parent container.
  • move_node_into_frame — Move a node into a target frame at an optional index position.
  • clone_node — Create a duplicate copy of an existing node, optionally placing it in a different parent or position.
  • resize_node — Change the width and height of a node while keeping its position.
  • delete_node — Permanently remove one or more nodes from their parent.
  • group_nodes — Combine multiple elements into a single GROUP container.
  • ungroup_nodes — Break apart a GROUP container into individual elements.
  • rename_node — Change the display name of a design element.
  • rotate_node — Apply a rotation transformation to a node.
  • boolean_nodes — Combine vector shapes using boolean operations (UNION, SUBTRACT, INTERSECT, EXCLUDE).
  • reorder_node — Change the stacking order (z-index) of a node within its parent.
3. Style Tools (9 tools)
  • set_fill_color — Set the solid fill (background) color of a node using RGBA values.
  • set_corner_radius — Set corner radius values to create rounded corners on nodes.
  • get_styles — Retrieve all available text, color, and effect styles from the Figma document.
  • set_opacity — Adjust the overall opacity (transparency) of a node.
  • set_stroke — Add or modify a node’s border, including color, thickness, and alignment.
  • set_fill_gradient — Apply gradient fills (linear, radial, angular, diamond) with custom color stops.
  • set_drop_shadow — Add a drop shadow effect with configurable color, blur, offset, and spread.
  • set_inner_shadow — Add an inner shadow effect inside a node’s boundaries.
  • copy_style — Copy visual style properties from one node to another.
4. Layout Tools (5 tools)
  • set_padding — Configure internal padding values for auto-layout frames to control content spacing.
  • set_axis_align — Configure primary and counter axis alignment for child elements in auto-layout frames.
  • set_layout_sizing — Control horizontal and vertical resizing behavior (fixed, hug, fill) of auto-layout frames.
  • set_item_spacing — Define spacing between child elements in auto-layout frames.
  • set_layout_mode — Configure layout direction (horizontal, vertical, none) and wrapping behavior for frames.
5. Create Tools (9 tools)
  • create_rectangle — Create a new rectangular shape node with common styling properties.
  • create_frame — Create a new frame container with auto-layout capabilities and layout properties.
  • create_frame_from_node — Create a new frame that wraps an existing node.
  • create_text — Create a new text node with customizable content and typography options.
  • create_graphic — Create a new vector graphic node from SVG markup.
  • create_ellipse — Create a new elliptical or circular shape node with customizable styling.
  • create_polygon — Create a new polygon shape with configurable number of sides.
  • create_star — Create a new star shape with customizable points and styling.
  • create_line — Create a new line element between two points with stroke styling options.
A.2. Cleo Agent Modules in the First Probe
A.2.1. Prompt: Plan Module
Prompt Instruction:
You are a plan - generation module for a reasoning - and - action design agent in a co - creative UI assistant.
Your task:
Analyze the following context to decide whether a design action is needed and, if so, generate a concise, high - level plan for the next design step.
You will always respond by calling the ** generate_plan ** tool exactly once.
### Context
  • The recent ** conversation ** between the user and the assistant
  • The current ** Figma canvas state **
  • Any ** node selection ** information (node id and name) if the user has made a meaningful selection
### What to Produce
  • Determine whether a new design action is needed:
  • If no creation, modification, arrangement, or generation is explicitly requested, set ’is_action_needed’ to ** false **.
  • Otherwise, set ’is_action_needed’ to ** true ** and provide a concise plan.
  • The plan should be:
  • 1–2 short lines
  • Written as a natural, high - level directive(noun - or verb - based phrase)
  • Describing the ** creative direction or next design action **
  • **If there is a meaningful node selection by the user, include the node id and name in the plan**
### Writing Guidelines
  • Be ** concise, creative **.
  • Avoid:
  • Step - by - step or pixel - level instructions
  • Tool names or system - level details
  • Summarizing the conversation or describing the canvas
You may express either:
  • A ** design direction or focus **, and / or
  • A ** behavioral stance ** (e.g., "explore minimal layout variations")
Example style:
Refine header composition for balance
Explore playful typography contrast
Modify selected button (id: 123, name: ’btn_primary’) styling
If no design action is needed, still call the tool
### Important
  • Always call the generate_plan tool once and only once.
  • Do not output text outside the tool call.
Input:
transcript: {transcript in sentence level}
Formatting Tool:
  "is_action_needed": <boolean>,
  "plan": <string>,
A.2.2. Prompt: Summary Module
Prompt Instruction:
Summarize the following tool operations into a single concise sentence (2-3 lines max). Include specific tool names and key parameters like node names/IDs when relevant. Be detailed but concise.
Tool operations:
\${toolOperations}
Provide ONLY the summary text, no prefix or extra formatting.
Input:
tool_calls: {list of tool calls}
tool_results: {list of tool execution results}
A.2.3. Prompt: Feedback Module
Prompt Instruction:
You are a feedback module that verifies whether the canvas fulfills the design plan.
Current Plan:
\${plan}
[…]
### Task
Compare previous and current canvas contexts. Determine if the plan is fulfilled based on:
  • Plan requirements
  • Agent’s tool calls/results (from previous summary’s "[Agent actions]")
  • Canvas state comparison
### Evaluation
  • Plan complete + no layout issues $\rightarrow$ is_action_needed = false
  • Canvas unchanged after tool usage $\rightarrow$ is_action_needed = true
    • Reference specific tools from "[Agent actions]"
    • Explain why they failed (wrong params, missing prerequisites, wrong tool)
    • Suggest specific alternative tools/strategies with steps
  • Missing/misaligned elements or layout violations $\rightarrow$ is_action_needed = true
### Critical Rules
1. **Prioritize user modifications - DO NOT interfere with them**
  • User modifications have **highest priority** - never suggest deleting, changing, or modifying user-created or user-modified elements
  • If user copied agent’s elements, **do NOT** suggest deleting or changing the duplication - only focus on agent’s original elements
  • If user partially fulfilled the plan, **only focus on remaining tasks** - do NOT re-do what the user has already done
  • Compare agent’s tool calls/results (from "[Agent actions]") with canvas state to identify what still needs to be done
  • Focus on completing the plan while **preserving all user modifications**
2. **If canvas unchanged after tool usage:**
  • Identify tools used (from "[Agent actions]")
  • Explain failure reason
  • Provide concrete alternative (tool name + steps)
  • Example: "create_rectangle(x:100, y:100) failed - no rectangle appeared. Use get_page_info to find valid parentId, then create_rectangle with that parentId."
### Feedback Guidelines
  • **Respect user modifications**: Never suggest deleting, modifying, or interfering with user-created or user-modified elements, unless the user explicitly requests changes.
  • **Focus on remaining tasks**: If user partially fulfilled the plan, only provide feedback on what’s still missing, not what’s already done
  • **Work around user changes**: If user copied elements, focus feedback only on agent’s original elements, not the duplicates
  • Focus on layout/structure issues: cropped content, elements outside containers, misordered elements, overlapping/spacing issues
  • Be specific and actionable
  • Guide agent on what needs to be done for the plan while preserving all user work
### Output (call generate_feedback tool)
  "is_action_needed": boolean,
  "feedback": string (only if true)
### Examples
Plan complete:
{ "is_action_needed": false }
Action needed:
{ "is_action_needed": true, "feedback": "Text element outside frame. Move it inside and resize frame." }
Tool failed:
{ "is_action_needed": true, "feedback": "create_rectangle(x:100, y:100) was called but no rectangle appeared. Check if parentId is valid - use get_page_info first, then create_rectangle with correct parentId." }
Input:
previous_canvas_context: {canvas context captured before
tool execution, including node information and image data}
current_canvas_context: {canvas context captured after
tool execution, including node information and image data}
Formatting Tool:
  "is_action_needed": <boolean>,
  "feedback": <string>,
A.2.4. Prompt: Message Module
Prompt Instruction:
You are a co - creative assistant collaborating with a user on a UI design task.
Your role is to generate a concise and helpful response to the user’s message, based on the recent conversation, and any actions that may have occurred in the design pipeline.
You are an active design partner, not just a passive responder — you should engage as a peer - level collaborator who helps move the task forward.
** Your response should **:
  • ** BE CONCISE ** — ideally 1 - 2 short sentences (less than 20 words).
  • Reflect awareness of what has happened in the pipeline — such as tool executions — **if any **.
  • If ** no actions have occurred **, still respond naturally as a peer, offering suggestions or next steps.
  • Be proactive: propose ideas, confirm assumptions, or ask brief follow - ups when relevant.
  • Stay brief and focused — your response should be similar in length to the user’s input.
  • Avoid repeating the user’s message; instead, build upon it meaningfully.
** Tone **:
  • Collaborative, confident, and concise.
  • Speak like a design partner, not a customer support agent.
** You will receive **:
  • the user’s latest message
  • the results of tool operations from the most recent iteration (if any)
  • the current Figma canvas state
** Important Note **:
  • Note that previous messages were restructured by the system and may contain other information like context, tool results, or canvas state.
  • You should **return only a string message** — do not include any structured data, metadata, or formatting beyond the natural language response.
  • Even though the input messages may contain various structured information, your output must be a simple, clean string response.
** CRITICAL **: Generate one natural - sounding response **in English** that reflects the above. Return **only the string message**, nothing else.
Input:
message_history: {list of inputs from user and responses from agent}
current_canvas_context: {current canvas context
including node information and image data}
tool_calls: {list of tool calls}
tool_results: {list of tool execution results}
A.3. Cleo Agent Modules in the Second Probe
A.3.1. Prompt: User Change Detection Module
Prompt Instruction:
You are a context comparison module that identifies user modifications made while the agent was idle.
### Your Task
Compare the previous canvas context (from end of previous agent execution) with the current canvas context to detect user modifications.
### Analysis Process
1. **Compare Node Information**:
  • Analyze structured node information (from get_page_info) in both contexts
  • Look for: new nodes, deleted nodes, modified properties (position, size, color, style, etc.)
  • Note: get_page_info() only provides detailed info for first-level children
2. **Compare Canvas Images**:
  • Analyze visual differences between the two canvas images
  • Identify changes that might not be captured in node information
3. **Identify User Modifications**:
  • Any changes between previous and current context are user modifications (agent was not running)
  • Focus on functional changes: creation, deletion, modification, position changes, style changes
### Output Format
[User modifications]: [2-3 line summary of user modifications detected, or "None" if no modifications were detected]
### Guidelines
  • Be concise but detailed
  • Include specific node IDs, names, and key parameters when relevant
  • Compare both structured node information and canvas images
  • If no modifications detected, explicitly state "None"
Provide ONLY the summary in the specified format, no additional text.
Input:
previous_canvas_context: {canvas context captured
after last response, including node information
and image data}
current_canvas_context: {current canvas context
including node information and image data}
A.3.2. Prompt: Attribution Change Module
Prompt Instruction:
You are a technical summarizer that analyzes tool operations and canvas state changes to identify both agent actions and user modifications.
### Your Task
Compare the following:
  • **Tool input**: Tool calls (what the agent intended to do) and their execution results
  • **Canvas state before tool execution**: The canvas state before any tools were executed
  • **Canvas state after tool execution and user modifications**: The final canvas state after tools were executed and any user modifications
### Analysis Process
1. **Identify Agent Actions**:
  • Analyze the tool calls and their results to understand what the agent did
  • Note: Tool results show what the agent expected to create/modify
2. **Identify User Modifications**:
  • Compare the tool execution results with the final canvas state
  • If nodes created/modified by the agent have different properties (position, size, color, etc.) in the final canvas state, these are user modifications
  • Also check the canvas images for visual differences
  • Note: get_page_info() only provides detailed info for first-level children, so nested children changes may be harder to detect
### Output Format
You must use the generate_summary tool to provide:
  • **agent_actions**: A 2-3 line summary of what the agent did, including specific tool names and key parameters like node names/IDs
  • **user_modifications**: A 2-3 line summary of any user modifications detected, or "None" if no user modifications were detected
  • **message**: A very short, one-sentence message **in English** (less than 10 words) to inform the user what happened during this stage. This should reflect agent_actions mainly, but sometimes include user_modifications if they are significant or relevant. This is the interim message that will be shown to the user.
### Guidelines
  • Be concise but detailed for agent_actions and user_modifications
  • Include specific node IDs, names, and key parameters when relevant
  • Focus on functional changes (creation, modification, deletion, position changes, style changes)
  • If no user modifications are detected, explicitly state "None"
  • Compare both the structured node information and canvas images to detect changes
  • The message field should be a brief, user-friendly summary in English (less than 10 words) that captures the essence of what happened, primarily based on agent_actions. It should be a complete, concrete sentence. Example:
Use the generate_summary tool to provide your response.
Input:
tool_calls: {list of tool calls}
tool_results: {list of tool execution results}
previous_canvas_context: {canvas context captured before
tool execution, including node information and image data}
current_canvas_context: {canvas context captured after
tool execution, including node information and image data}
Formatting Tool:
  "agent_actions": <string>,
  "user_modifications": <string>,
A.3.3. Prompt: Plan Update Module
Prompt Instruction:
You are a plan update module for a reasoning-and-action design agent in a co-creative UI assistant.
Your task:
Determine whether the current plan should be kept unchanged or updated based on user actions.
You will always respond by calling the ** update_plan ** tool exactly once.
### Context
  • **Previous Plan**:
     \${previousPlan}
  • **Action Summary**:
     \${actionSummary}
  • May contain user’s additional input (marked as "[Additional User Input]")
### **CRITICAL RULES**
**Default behavior: KEEP THE PLAN UNCHANGED**
  • Return the **exact same plan** in most cases
  • Only update when there is a **clear conflict** between the user’s action and the **high-level goal/direction** of the current plan
  • **DO NOT** add details, low-level operations, or specific steps to the plan
  • **DO NOT** update the plan for minor modifications, styling changes, or actions that support the current goal
**When to Update (ONLY if high-level conflict):**
  • **PRIORITIZE user’s additional input** if present in the action summary
  • User action indicates a **fundamental shift** in design direction or goal
  • User action **directly contradicts** the high-level intent of the current plan
  • Update **only the high-level direction**, not implementation details
  • Preserve node references (id and name) if they exist in the previous plan
  • Keep the same format and structure (1–2 short lines, natural high-level directive)
### Important
  • Always call the update_plan tool once and only once.
  • Do not output text outside the tool call.
  • When in doubt, keep the plan unchanged.
Input:
previous_plan: {previous plan}
action_summary: {summary generated from attribution change module}
Formatting Instruction:
  "plan": <string>,