### 07 Apr 2026 v4

# “When to Hand Off, When to Work Together”: Expanding Human-Agent Co-Creative Collaboration through Concurrent Interaction

Abstract.
As agents move into shared workspaces and their execution becomes visible, human-agent collaboration faces a fundamental shift from sequential delegation to concurrent co-creation. This raises a new coordination problem: what interaction patterns emerge, and what agent capabilities are required to support them? Study 1 (N=10) revealed that process visibility naturally prompted concurrent intervention, but exposed a critical capability gap: agents lacked the collaborative context awareness needed to distinguish user feedback from independent parallel work. This motivated Cleo, a design probe that embodies this capability, interpreting concurrent user actions as feedback or independent work and adapting execution accordingly. Study 2 (N=10) analyzed 214 turn-level interactions, identifying a taxonomy of five action patterns and ten codes, along with six triggers and four enabling factors explaining when and why users shift between collaboration modes. Concurrent interaction appeared in 31.8% of turns. We present a decision model, design implications, and an annotated dataset, positioning concurrent interaction as what makes delegation work better.
Human-Agent Interaction, Interaction Pattern Analysis, Concurrent Interaction, Collaborative Agent, Creative Workflow
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Interactive systems and tools
1. Introduction
AI agents are increasingly capable of executing complex tasks autonomously, such as design (Make, ; Stitch, ; Lovable, ), programming (Sapkota et al., 2025), and web navigation (Jiang and Xia, 2025). As agents move into shared workspaces and their execution becomes visible through emerging tool-calling infrastructures (e.g., Model Context Protocol (Anthropic, 2024)), human-agent collaboration faces a fundamental shift: from sequential delegation, where users specify goals and review outputs, to collaborative co-creation, where users and agents work simultaneously in the same space. This shift raises a new coordination problem: what interaction patterns emerge when humans and agents share a workspace, and what agent capabilities are required to support them?
To explore this, we built a process-transparent agent in Figma that executes design tasks through direct tool operations (e.g., creating frames, applying shadows), helping designers track the agent and coordinate accordingly (Gutwin and Greenberg, 2002; Gutwin et al., 1996; Robertson, 2002). In Study 1 with 10 designers, process visibility revealed how the coordination problem manifests. Designers naturally began intervening mid-execution, interleaving their own work with the agent’s and demonstrating intent through direct manipulation, creating a need for the agent to interpret and respond to these concurrent actions (Figure 1). However, the agent could not interpret these concurrent actions, lacking the collaborative context awareness needed to distinguish feedback to incorporate from independent work to leave alone.
Figure 1. An example of collaborative friction in human-agent collaboration within a shared workspace.
The figure illustrates a conflict scenario in human-agent collaboration on a shared canvas. In the initial state (top), the agent is executing a two-column layout while narrating its progress, and the user concurrently proposes a three-column layout. The circular arrows indicate both parties acting simultaneously on the same workspace. In the resulting state (bottom), the user’s three-column edit is visible on the canvas, but the agent — unaware of the user’s intent — treats it as an error conflicting with its original plan and reverts to the two-column layout. The red spiky border highlights the moment of conflict.
Based on this finding, we developed Cleo (Collaborative Linked Executive Operator), a second design probe that embodies collaborative context awareness by continuously tracking user actions within the workspace, interpreting whether these actions should update the agent’s task planning or be considered as independent work, and adapting the agent’s ongoing execution accordingly (Horvitz, 1999; Pellegrinelli et al., 2016).
With Cleo, we conducted a two-day exploratory study (Study 2, N=10) with stimulated recall interviews (Dempsey, 2010; Calderhead, 1981), analyzing 214 turn-level interactions to understand when and why users choose different collaboration modes. We identified a taxonomy of five action pattern categories and ten codes capturing how users engage with agents during execution (hands-off, observational, concurrent, directive, terminating), six triggers prompting intervention (e.g., idea spark, task mismatch), and four enabling factors shaping those choices (e.g., mental model, task priority). Concurrent work appeared in 31.8% of turns, with dynamic shifts driven by evolving priorities. We also synthesized these findings into a decision model (Figure 6) with six interaction loops, providing a grounded framework for understanding when users expect intervention support, when they prefer autonomous execution, and how agents can anticipate and adapt to these fluid transitions.
Drawing from these findings, we derive design implications for adaptive collaborative agents, examine concurrent interactions as behavioral signals for workstyle learning, and discuss how these principles extend beyond creative domains. Our findings position concurrent interaction not as a replacement for delegation, but as what makes delegation work better. By building mutual visibility and understanding that allows agents to learn how users work, full delegation becomes not a leap of faith, but a natural extension of an already-established collaboration. We also release our codes with a dataset of 214 interaction turns to support future research on human-agent co-creative collaboration by simulation. Our contributions include:
  • A taxonomy of user action patterns in human-agent interaction, comprising five categories and ten codes characterizing how users engage with agents during execution.
  • A decision model with six interaction loops explaining how users navigate between delegation, direction, and concurrent work based on six triggers and four enabling factors.
  • Design implications for adaptive collaborative agents, covering process visibility adaptation, behavioral signal interpretation, and proactive execution management.
  • A dataset of 214 human-agent co-creative interaction turns, uniquely capturing concurrent interaction logs (when and how users work alongside agents)
2. Related Work
2.1. Dynamics of Flexible Collaboration: From Turn-taking to Synchrony
Human collaboration is a dynamic spectrum of interaction modes (Tang, 1991). While turn-taking reduces cognitive load and conflict in complex tasks through sequential actions (Kirschner et al., 2009; Sacks et al., 1974), creative domains often demand synchronous collaboration (Nam and Sakong, 2009). Creative work involves the co-evolution of problem and solution, where design ideas are refined through immediate visual feedback and spontaneous mutual inspiration (Maher and Poon, 1996; Ishii et al., 1993). Therefore, partners often bump in for immediate adjustments or perform simultaneous edits to interleave workflows, ensuring that the evolving design remains a shared mental model (Dourish and Bellotti, 1992; Tang, 1991).
Modern tools like Figma and Google Slides support this fluidity by fostering workspace awareness. This is enabled through visual cues like telepointers (moving cursors) and object highlighting, allowing users to anticipate moves and intervene without explicit communication (Gutwin and Greenberg, 2002; Gutwin et al., 1996). However, most co-creative AI systems remain confined to a rigid wait-your-turn paradigm, largely limiting human-AI interaction to a prompt-and-output cycle, lacking the mutual awareness essential for true creative partnership. Moreover, working with generative AI has been shown to reduce divergent thinking (Anderson et al., 2024), underscoring that how humans and AI collaborate matters as much as whether they do.
2.2. Paradigm Shifts in Human-Agent Collaboration
AI is evolving from a command-based tool into a proactive “co-worker” that actively participates within shared workspaces (Horvitz, 1999). This shift has been driven by frameworks enabling models to leverage diverse tools (Schick et al., 2023) and interleave reasoning with actions (Yao et al., 2023), enabling agents for programming (by Anthropic, ; Jimenez et al., 2023; Sapkota et al., 2025), web automation (AI, ; Deng et al., 2023; Jiang and Xia, 2025), visual design (Make, ), research (Gottweis et al., 2025; Feng et al., 2024), and data analysis (Hong et al., 2025). As these agents expose their reasoning and intermediate actions, this affords greater interpretability by allowing users to monitor agent behavior in real time (Amershi et al., 2019; Prasongpongchai et al., 2025).
Yet most collaboration between human users and AI agents remains rooted in turn-taking. This limits users’ ability to correct agents mid-execution (Amershi et al., 2019) to prevent misalignments as the task progresses. In this work, we explore the potential of concurrent human-agent collaboration not as a replacement for turn-taking, but as an additional mode—enabling agents to move beyond delegated executors to responsive partners capable of accommodating spontaneous interventions within a shared workspace. This reflects a broader recognition that full delegation—optimizing purely for efficiency—risks devaluing the creative process itself (Almeda et al., 2025).
2.3. Design Agents: From Workflow Automation to Process-Visible Collaboration
Existing agents for supporting the design process can be categorized into three distinct approaches based on how they structure their tasks and interact with users. The first category is subtask design agents, which support specific stages within an existing design workflow by automating specific and tedious subtasks, adding details (Lawton et al., 2023; Karimi et al., 2020; Zhou et al., 2024) or generating alternatives (Make, ), while the overall workflow remains human-driven. The second, process-structured agents, operate through predefined node-based workflows (Hao et al., 2025; Wang et al., 2025; Ding et al., 2023; You et al., 2025), supporting guided co-creation, but constrained by predefined logical nodes. The third, output-centric agents, expose only a textual reasoning trace before delivering final results (Stitch, ; Lovart, ; Khurana et al., 2025).
Across all three, the agent’s execution remains hidden from the user’s workspace: users interact only through requests and final outputs, with no visibility into work in progress. While agents in other domains (e.g., programming) increasingly expose execution steps through direct tool operations, design tools have yet to explore how “live” tool-use—such as visible cursor movements and object manipulation—could reshape human-agent collaboration. This suggests an open question of whether process visibility could enable more diverse collaboration modes beyond simple turn-taking.
3. Study 1: Initial Technology Probe Study
To explore how process visibility creates opportunities for co-creative interactions and reveals the coordination challenges it entails in creative workflow, we implemented an initial design probe and conducted an exploratory study with 10 designers.
3.1. First Probe System
3.1.1. Three Basic Components for Collaborative Agent Design
Existing design agents are largely opaque—they accept instructions and return text reasoning (or high-level sub-goals) and final outputs, giving users little visibility into execution. To make agent execution observable and collaborative, we grounded the probe’s design consideration in three core components drawn from human-human collaboration literature, where these principles have been shown to support coordination, mutual understanding, and fluid collaboration, reasoning that they could equally scaffold more natural human-agent collaboration: (1) Interaction Modality for peer-like communication, (2) Process Transparency for observable agent work, and (3) Presence Awareness for mutual attention visualization.
(1) Interaction Modality: Voice Interaction with Flexible Object Referencing. The first probe uses voice as the primary interaction channel, enabling natural turn-taking, real-time intention sharing, and a sense of horizontal partnership (Hutchins, 1995; Reicherts et al., 2022; Bickmore and Picard, 2005). Users can also reference canvas objects directly during speech, enabling fluid deictic communication (e.g., “make this larger” while clicking).
(2) Process Transparency: Sharing Implementation Workflow. The first probe system exposes its design process in real-time through step-by-step execution, so users observe not just what the agent produces but how it progresses, establishing common ground around both goals and process (Robertson, 2002; Cila, 2022; Hill et al., 1992).
(3) Presence Awareness: Visualizing Attention in the Shared Workspace Following Gutwin’s workspace awareness framework (Gutwin and Greenberg, 2002), the agent’s position reveals its current focus, enabling users to anticipate upcoming actions and coordinate accordingly (Dourish and Bellotti, 1992; Fussell et al., 2000).
3.1.2. System Design
Figure 2. Interaction flow of the first probe system. (a) The user sends a voice message (b) while selecting canvas elements for context. (c) The agent generates an action plan based on the voice transcription with selections and iterates through ReAct cycles. (d-f) Each iteration’s progress is reflected immediately on the canvas as tools execute during the act stage.
This figure illustrates the interaction flow of the first probe system through an example scenario in which a user requests color variations for a selected section. The user delivers a voice request while selecting relevant canvas elements, allowing the system to capture both the verbal intent and the visual context. Based on this input, the agent generates an action plan and iteratively progresses through cycles of reasoning, tool execution, and feedback evaluation. During each iteration, the agent executes design tools directly on the canvas and provides interim messages, enabling users to observe incremental changes and intervene if needed before the final output is produced.
Interface and Interaction Scenario. The first probe is implemented as a Figma plugin where users combine voice input with canvas element selection to provide canvas context-based instructions to the agent (Figure 2-a-b). The agent responds through iterative tool execution—reasoning, acting, and self-evaluating in cycles—with each step reflected immediately on the canvas via a floating agent character alongside interim messages. (Figure 2-c-f).
Agent Design and Pipeline. The first probe agent is built on the ReAct structure (Yao et al., 2023), operating through a three-step iteration cycle: (1) reason, (2) act, and (3) feedback (Figure 2-c; Appendix A.4). At the reasoning stage, the agent selects from 38 Figma operations (e.g., create rectangle) and generates an interim progress message, which is executed on the canvas alongside the agent character. After each cycle, the summary module generates a text description of what occurred, while the feedback module compares the canvas state before and after execution to evaluate plan completion. If the plan is fulfilled, the agent terminates and delivers a final response; otherwise, it begins a new iteration, up to a maximum of 10 cycles .
We use Claude Sonnet 4.5 (Anthropic, 2026b) as the main reasoning model and Claude Haiku 4.5 (Anthropic, 2026a) for all other modules. Full prompt instructions, including Figma design guidelines (adapted from (Jeong et al., 2026)) and the complete set of 38 Figma operation tools (See Appendix A.1.2).
3.2. Study Design
Study 1 investigates two questions: (1) How does process transparency—observing the agent’s step-by-step workflow on the canvas—affect users’ collaboration experience and enable a new type of collaboration? (2) What design considerations are needed for better human-agent collaboration in co-creative workflow? We observed 10 professional UI/UX designers collaborating with our design probe across different task types to examine how process transparency shapes their understanding and collaboration strategies.
3.2.1. Participants
We recruited 10 UI/UX designers (D1-D10, 3 male, 7 female, ages 24-36, mean=28.30) with at least 1 year of professional design experience. We screened for high Figma proficiency (e.g., daily users) and prior experience with diverse design agents (e.g., Figma Make (Make, ) or Lovart (Lovart, )). Participants were compensated 40,000 KRW for approximately 1.5 hours of participation.
3.2.2. Procedure
After introducing different types of design agents (e.g., OptiMuse (Zhou et al., 2024) and Figma Make (Make, )) and a brief tutorial on the probe’s three interaction components, participants completed two design tasks of varying scope to observe how collaboration strategies differ across task complexity: (1) a fine-grained task designing a single webpage section (e.g., hero or project gallery) with given materials (15 mins), and (2) a coarse-grained task creating an entire webpage (e.g., startup landing page) across multiple sections (25 mins). We concluded with a 30-minute semi-structured interview exploring how process transparency shaped their collaboration strategies and interaction with the probe.
3.3. Findings
3.3.1. Process Transparency: Real-Time Understanding and Intervention Opportunities
Observing the agent’s step-by-step execution fundamentally transformed how designers experienced collaboration. All designers reported a strong sense of “working with the agent” rather than delegating to then reviewing it. Unlike traditional agents, where users assign tasks and wait for final outputs, process transparency enables concurrent work: participants understand what the agent is doing, how it will end up, and plan or redirect their own tasks based on this understanding.
This transparency relieved the “loss of agency” problem common in human-agent interaction (Schömbs et al., 2026). Participants contrasted our probe with general design agents, where agents take over the workspace, leaving users waiting passively and fearing disruption if they intervene. With those agents, users must read lengthy reasoning texts to understand their plans or reasoning—a cognitively demanding process. In contrast, seeing intermediate behaviors with output on the canvas made the agent’s intentions intuitively understandable. D4 commented “I don’t need to read paragraph. I just see it creating a button, and I know what’s happening.” This understanding of how the agent works prevented over-dependence on opaque automation, preserving user agency throughout the process.
Most critically, step-by-step observation enabled early evaluation and proactive intervention. Participants could anticipate the agent’s next steps (D1-D10), identify the agent’s misinterpretation on the given task (D1, D3, and D5-D7), and recognize their own unclear instructions (D1-D2, D4, and D8), and assess the agent’s overall workflow direction (D2, D6, and D8-D10)—all before the agent finished. Importantly, this understanding enabled participants to plan their own flexible parallel work: while the agent executed, participants gathered reference images, evaluated overall design directions, or prepared content for subsequent sections, rather than passively waiting. Specifically, D7 stated: “I saw it working on the hero section with a certain direction, so I started looking for references related to that direction for the next section instead of just watching.” This predictability helped participants manage workflows more effectively, deciding when to intervene, to prepare subsequent tasks, and to let the agent continue autonomously. Process transparency thus shifted collaboration from reactive correction of final outputs to proactive coordination during execution.
3.3.2. Emerging Expectations: Flexible Concurrent Collaboration and Context Awareness
Agent’s process transparency fundamentally changed the users’ expectations: they wanted to engage directly with the agent’s work-in-progress through concurrent intervention rather than waiting for completion. For example, participants wanted to appropriate intermediate outputs for their own designs, assist the agent by working on related elements simultaneously, or demonstrate desired changes through direct editing. Most of the participants (D1-D3 and D6-D10) explained that observing the process made them more aware of how they wanted things done, creating natural moments to intervene and collaborate. D3 noted: “I could see exactly where it needed my input—I wanted to just jump in and show it rather than explain this verbally.”
However, the first probe’s inability to distinguish between user and agent actions in the canvas created conflicts. Because the agent tracked overall canvas changes without separating user modifications, it sometimes reverted or misinterpreted concurrent user work. Participants who experienced this avoided intervening during execution, relegating their collaboration ideas to post-task interviews. Yet even participants who avoided intervention reported that their reluctance to interrupt agents stems from prior experiences with other agent systems (D2, D4-D7, D10)—fearing that mid-execution changes would derail the agent’s workflow. Process transparency challenged this assumption; designers recognized opportune moments for intervention but lacked system support.
Participants envisioned real-time interleaved workflows where users could intervene during the agent’s ongoing execution: appropriating agent-created elements into their own designs while the agent continued working, handing off their work-in-progress for the agent to build upon mid-execution, or collaboratively editing the same artifact simultaneously. Critically, they expected agents “to understand why they intervened”—whether corrections indicated misunderstanding, appropriations showed approval, or demonstrations communicated desired style (D1, D3-D5, D7-D8). This requires the agent’s collaborative context awareness: distinguishing feedback on agent work from independent user work and recognizing when users want the agent to adapt versus continue autonomously. Process transparency thus revealed not just new collaboration opportunities, but the need for context-aware agents that support—rather than conflict with—real-time interleaved human actions.
4. Study 2: Two-Day Technology Probe Study with Stimulated Recall Interviews
While Study 1 revealed the coordination problem that arises when agents lack collaborative context awareness, it remained unclear what interaction patterns emerge in the collaboration and what enables users to shift between delegation, direction, and concurrent work. To investigate this, we conducted a two-day technology probe study (Hutchinson et al., 2003) with stimulated recall interviews (Dempsey, 2010; Calderhead, 1981) observing experienced creative professionals collaborate with the probe.
4.1. Second Probe System: Cleo
(Collaborative Linked Executive Operator)
4.1.1. Understanding Collaborative Context for Flexible Human-Agent Co-Creation
Study 1 revealed that process transparency prompts users to interleave their actions with the agent’s ongoing execution. The key challenge is collaborative context awareness—agents must distinguish whether concurrent user actions represent feedback requiring adaptation, independent parallel work to be preserved, or co-editing to be incorporated into ongoing execution.
Therefore, we developed Cleo (Collaborative Linked Executive Operator), which extends the first probe by incorporating collaborative context awareness as its core principle, enabling the agent to recognize and adapt to users’ direct manipulation during execution.
4.1.2. Cleo Design and Improvements from Design Probe 1
Figure 3. Flexible co-creation scenario with Cleo. (a) The user invokes Cleo by calling its name. (b) While Cleo is acting on the canvas, the user can (c) work concurrently on the same task, or (d) intervene directly with additional instructions at any time.
This figure illustrates a flexible co-creation scenario enabled by Cleo, highlighting how users can dynamically shift their level of involvement during an ongoing task. After invoking the agent to create a portfolio header, the user observes the agent’s initial output on the canvas. In subsequent iterations, the user may abort the agent’s operation to work independently, make concurrent edits directly on the canvas while the agent is active, or provide additional verbal instructions to redirect the agent’s behavior. The sequence demonstrates how Cleo supports fluid transitions between hands-off delegation, concurrent collaboration, and direct intervention within a single task.
Interaction Improvements. Building on the first probe system, Cleo supports three additional interaction types (Figure 3). First, users can abort the agent’s operation at any time via the termination button on the plugin. Second, users can work concurrently with the agent on the same elements, with the agent tracking and prioritizing user operations to align its plan accordingly (Figure 3-c). Third, users can provide additional voice instructions at any time by calling the agent’s name while it is working, which the agent incorporates in the next iteration (Figure 3-d).
Technical Improvements. To support these, we replaced the original Summary Module with three new modules (Figure 4). The User Change Detection Module analyzes user actions that occurred while the agent was inactive, maintaining awareness of the overall workflow. The Attribution Change Module distinguishes between agent and user actions during execution, producing separate summaries for each. The Plan Update Module then updates the current plan only when user actions conflict with or indicate new directions from the existing plan, before initiating the next iteration cycle.
Figure 4. Agent structure of Cleo. Three modules have been updated from the first probe pipeline (Appendix A.4): User Change Detection Module, Attribution Change Module, and Plan Update Module.
This figure presents the updated agent architecture of Cleo, highlighting three additional modules introduced beyond the first probe pipeline to support user interventions during agent execution. Before planning, the User Change Detection Module analyzes user actions that occurred while the agent was inactive and incorporates them into the chat history and canvas context. During execution, the Attribution Change Module distinguishes between agent-generated changes and user modifications by processing tool call results together with the current canvas state. Based on this attribution, the Plan Update Module selectively revises the agent’s plan when user actions conflict with the current plan or indicate new directions. These updates are integrated into the existing ReAct-based loop of tool execution, feedback evaluation, and plan fulfillment, enabling adaptive planning in response to ongoing user interaction.
Figure 5. How Cleo achieves context awareness. (a) The user invokes the agent. (b) It recalls past observations and detects user updates against the current state. (c) It monitors concurrent user actions while executing its plan, prioritizes interventions, and updates its plan as needed.
The figure demonstrates how the system achieves context awareness through three components. In (a), the user invokes the agent by issuing a voice command to add icons to the footer of a portfolio canvas. In (b), the User Change Detection Module compares the agent’s previous observation — a simple footer with email and contact fields — against the current canvas state, detecting that the user has updated the footer content and added contact information. In (c), while the agent is actively creating icons on the canvas, the Attribution Change Module detects that the user concurrently changed the layout alignment to vertical. This intervention triggers the Plan Update Module, allowing the agent to revise its plan accordingly.
4.2. Study Design
We conducted a two-day exploratory study with Cleo to understand what interaction patterns emerge when users collaborate with a context-aware agent, and what enables or triggers different collaboration modes. By combining interaction logs with stimulated recall interviews, we aimed to capture not just what users did, but why they made those choices in context.
On the first day, participants performed two design tasks (closed- and open-ended) in Figma, with interactions logged and screen-recorded. After the session, three authors visually represented the interaction logs by segmenting the entire task process into agent-activated and agent-deactivated periods, and labeling each segment with timestamped user and agent actions to create an easily reviewable format for the stimulated recall interview ( Appendix B.1-c-d). We gave a one-day gap between the task session and recall interview to avoid participant fatigue and for visualizing logs. On the second day, stimulated recall interviews were conducted where participants reviewed these visualized interaction logs.
4.2.1. Participants
We recruited 10 participants (4 males and 6 females; M_age = 30.1, SD_age = 3.9) who met two key criteria: (1) at least two years of professional design experience, and (2) active use of VLMs (e.g., ChatGPT (OpenAI, 2026) or Gemini (Google, 2026)) or design agents (e.g., Figma Make (Make, ), Stitch (Stitch, ), or Lovable (Lovable, )) in their creative workflow. Participants were compensated 120,000 KRW for approximately 3 hours of participation.
4.2.2. Procedure
First day. After an introduction, participants familiarized themselves with Cleo through a warm-up task (30 mins), followed by two main tasks: (1) a closed-ended task (30 mins) where participants designed an empty section of a web portfolio to match existing sections given design requirements and source materials, and (2) a open-ended task (40 mins) where participants selected one of three startup concepts and freely designed a webpage from scratch.
After the tasks, three authors processed the collected interaction logs by marking timestamps of user requests and agent task completion. Using these timestamps, each participant’s entire task process was segmented into agent-activated and agent-deactivated periods, then grouped by each user request. Segmented video clips were also prepared alongside these logs.
After these tasks concluded, three authors prepared visualizations of the interaction dynamics to facilitate the stimulated recall interview on the second day. The authors first marked the moments when participants sent requests to the agent and when the agent completed execution for each request (Appendix B.1-a). Using these two time points to segment the entire task workflow, the authors visualized the data (Appendix B.1-b-d), distinguishing between periods when the agent was actively working and when it was not. Alongside these visualizations, we prepared segmented video clips of the actual task recordings.
Second day. Participants reviewed their visualized interaction logs and corresponding video clips. For each segment, we inquired about the reasons and purpose behind their actions, and the collaboration contexts where such actions could be effective within 1 hour. A 20-minute in-depth interview followed by exploring overall collaboration experiences, the effectiveness of concurrent work, and what implicit information their actions conveyed to the agent.
4.2.3. Analysis Methods
Qualitative coding of the user’s action patterns. From the study, we collected 214 turn-level data from 10 participants across two tasks, where a turn means a complete agent execution in response to a user request comprising up to 10 iterations (Section 2), with an average of 21.4 turns per participant (min=12, max=33). Three authors conducted open coding (Corbin and Strauss, 1990; Khandkar, 2009) on what actions users performed while the agent was working, separating multiple actions within a single turn for individual coding. Two authors iteratively reviewed the data to consolidate overlapping actions into broader categories, after which a third author joined to refine and finalize until consensus was reached. As a result, we defined 5 categories and 10 codes (Table 1).
Distribution analysis of user action patterns. After coding, we annotated each of the 214 turns with the 5 categories, where each turn could be mapped to 1 to 5 categories (Appendix B.3), and calculated each category’s proportion across all 214 turns.
Qualitative coding of triggers and enabling factors of user actions. Next, we conducted qualitative coding on the stimulated recall interview data to understand when and why users chose different action categories, identifying two key pattern types: triggers that prompted transitions from observation to intervention, and enabling factors—situational conditions including user’s mental models, task priorities, and intervention preferences—that determined which specific action users adopted. Following the same methodology as above, with the addition of a fourth author who had not been involved in conducting the study sessions to provide an independent perspective and reduce potential bias.
Iterative decision modeling of human-agent co-creative collaboration. Building on the identified action patterns, triggers, and enabling factors, we constructed an initial decision model describing how user-agent collaboration unfolds from task initiation to completion, representing transitions between different user actions and the conditions under which these transitions occur. Through a constant comparative process (Glaser and Strauss, 2017), we iteratively validated and refined the model against all 214 turn-level processes, revising it whenever an observed turn sequence could not be adequately explained. This process was conducted by five authors, including an additional author uninvolved in the study sessions to provide an independent perspective, and continued until the model consistently accounted for all cases without contradiction.
4.3. Results
We present five user action patterns and their distribution across 214 turns, followed by six triggers and four enabling factors that explain when and why users chose each pattern, and synthesize these findings into a decision model of human-agent co-creative interaction.
4.3.1. User Action Patterns During Agent Execution
[…]
(H) Hands-off Full delegation Completely disengaging from the agent’s work to focus on other tasks. 70.09%
(O) Observational Observational monitoring Watching the agent’s execution without intervening. 68.69%
(T) Terminating Execution termination Stopping the agent’s work before completion. 8.88%
(D) Directive Instruction-based steering Providing verbal guidance to adjust the agent’s approach. 28.50%
Switching tasks Assigning a new task unrelated to the current execution.
(C) Concurrent Intermediate result appropriation Copying and using the agent’s partial outputs while it continues working. 31.78%
Artifact takeover Duplicating the agent’s work-in-progress to edit independently elsewhere.
In-situ co-editing Working simultaneously with the agent on the same subtask and artifact.
Opportunistic takeover Completing one of the agent’s pending subtasks while it handles another.
Demonstration-based steering Showing desired changes through direct editing rather than instructions.
Table 1. Five categories and ten codes identified through qualitative analysis of user action patterns. See the detailed figures of each interaction in Appendix 9
This table presents a taxonomy of user intervention patterns when collaborating with an AI agent. It categorizes ten distinct action codes into five high-level categories—Hands-off, Observational, Terminating, Directive, and Concurrent—each describing different levels and modes of user engagement, from full delegation to simultaneous co-editing.
Through the first qualitative coding process, we identified 5 categories and 10 action patterns users took while the agent was working (Table 1). Hereafter, each category is denoted by its initial: H Hands-off, O Observational, T Terminating, D Directive, and C Concurrent. These patterns can be characterized along four key dimensions: intervention modality (whether users intervene through verbal instructions or direct actions on the agent’s workflow), artifact coupling (whether the user and agent work on separate or overlapping artifacts), synchronicity (whether interactions occur synchronously or asynchronously with agent execution), and task autonomy (who has primary control over the task execution).
Notably, the C category represents a novel set of action patterns enabled by the agent’s collaborative context-awareness (Appendix B.2). Unlike traditional design agents that block user intervention during execution (Make, ; Lovable, ) or coding agents where direct interventions are not reflected in the agent’s work (by Anthropic, ; by Anysphere, ), Cleo allows users to directly engage with the agent’s ongoing work without disrupting its workflow. This enables more precise communication of nuanced and situational intent—particularly for aesthetic decisions that are difficult to articulate verbally—as observing the agent’s step-by-step execution often surfaced latent preferences users had not previously considered. Many participants noted that direct manipulation was easier than verbal explanation for such fine-grained adjustments, prompting them to show the agent what they meant directly (P1-P3, P5, P7-P8, and P10).
Within the C category, In-situ co-editing and Opportunistic takeover both involve synchronous direct manipulation on overlapping artifacts, yet differ in task autonomy: the former creates shared control over the agent’s task (Agent → Shared) while the latter maintains the agent’s autonomy over its assigned subtask as the user completes a different pending subtask. Similarly, Artifact takeover and Intermediate result appropriation both shift to shared control but differ in scope: the former involves duplicating the agent’s entire work-in-progress to edit independently, while the latter involves copying only partial outputs.
The D category distinguishes between Instruction-based steering, where users provide verbal guidance while maintaining the agent’s autonomy, and Switching tasks, where users assign a new task while the agent is still executing the previous one.
Execution termination of T category shifts task control entirely to the user (Agent → User) by stopping the agent’s work completely.
Turns ranged from containing only one category to including all five categories (Appendix B.3). Among 214 turns, the most prevalent turn type was immediate hands-off (turns only include H), accounting for 31.3% of all turns, followed by observation-only turns (turns only include O) at 14.0%. When we calculated how each action category appeared across 214 turns, H appeared in 70.1% of turns, O in 68.7%, C in 31.8%, D in 28.5%, and T in 8.9% (Table 1’s % of Turns).
4.3.2. Triggers and Enabling Factors
Through the second qualitative coding process, we identified six triggers that lead users to perform actions in the C, D, and T categories (Table 2), along with four enabling factor categories that determine which action category users choose when these triggers occur (Table 3).
O
occurred in all turns except those containing only H, consistently preceding C, D, and T action categories. However, O does not always follow H—some turns showed users initially working hands-off H, then later shifting to O. The C, D, and T categories were triggered during this observation process. Thus, we define triggers as the reasons that lead users to perform one of these three action categories during the observation process.
[…]
Idea Spark from Agent’s Work-in-Progress $\checkmark$
Need for Early Outcome Visibility $\checkmark$
Readiness for Fine-grained Detailing $\checkmark$
Misaligned Task Interpretation $\checkmark$ $\checkmark$ $\checkmark$
Execution Quality Drop $\checkmark$ $\checkmark$ $\checkmark$
Emerging New Task for Agent $\checkmark$ $\checkmark$
Table 2. Six triggers that prompt user actions across the C, D, and T user action categories.
[…]
Has mental model $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
(a) Mental Model of Agent’s Task Capability No mental model $\checkmark$
User task significantly more important $\checkmark$
User task moderately more important / Similar importance $\checkmark$ $\checkmark$ $\checkmark$
(b) Task Importance: User vs. Agent No user task $\checkmark$ $\checkmark$ $\checkmark$
Verbal explanation easier $\checkmark$
Direct manipulation easier $\checkmark$
(c) User’s Preferred Intervention Modality Uncertain how to intervene $\checkmark$ No intervention
Task will succeed with direct collaboration $\checkmark$
Agent will understand verbal guidance $\checkmark$
(d) User’s Expectation of Agent’s Response to Intervention Agent is incapable of this task $\checkmark$ No intervention
Table 3. Four categories and eleven codes of enabling factors across action categories.
A total of six triggers were identified (Table 2). Three triggers exclusively led to Concurrent interactions. Idea Spark from Agent’s Work-in-Progress occurred when observing the agent’s intermediate outputs sparked new creative ideas that the user had not previously considered. Need for Early Outcome Visibility arose when users wanted to see or utilize the agent’s results sooner to plan subsequent tasks. Readiness for Fine-grained Detailing emerged when users recognized that the current work stage required detailed refinement aligned with their specific vision.
Two triggers could lead to any of the three action categories ( C, D, or T). Misaligned Task Interpretation happened when the agent pursued a valid but unintended interpretation of the task, executing in a direction the user did not intend. Execution Quality Drop was triggered when the agent’s performance fell below the user’s acceptable threshold.
Finally, Emerging New Task for Agent exclusively led to D or T, occurring when users conceived a new task to assign to the agent while observing the agent’s work.
We identified four enabling factor categories and eleven codes (Table 3). The first factor (Table 3-a) concerns whether the user has a mental model of the agent’s capability for the assigned task. All action categories except O were performed when a mental model existed, but O also emerged when no mental model was present. For instance, many participants closely monitored the agent’s work from start to finish when they lacked a mental model of its capabilities, and even when triggers occurred, they did not proceed to T, C, or D actions in order to first observe how well the agent could perform the given task.
The second factor (Table 3-b) concerns the user’s perceived importance of their current work relative to the agent’s task. When the user’s work requires full concentration, they move to H; when importance is similar or the user has no work to attend to, they remain available to observe and intervene, with C occurring when the user has no work and D when their work is of similar or slightly greater importance.
The third factor (Table 3-c) concerns the user’s preferred intervention modality. When verbal explanation is easier, users choose D actions (e.g., “Please modify this design based on this reference”). When direct manipulation is easier, users choose C actions, as observed in participants who preferred to directly adjust fine-grained details such as alignment spacing rather than describe them verbally (P1, P3, and P9). When users are uncertain about how to intervene, they choose to terminate the agent’s workflow T.
The final factor (Table 3-d) concerns the user’s expectation of which intervention approach will enable the agent to produce a successful outcome. When users believe direct manipulation alongside the agent will lead to successful completion, they choose C; when they expect verbal feedback alone will be sufficient, they choose D; and when they anticipate the agent cannot perform the task regardless of intervention, they choose T.
4.3.3. Decision model of Human-Agent Collaboration
Figure 6. Decision model of human-agent co-creative collaboration
This figure presents a flowchart-style decision model of human–agent co-creative collaboration. The diagram starts from an initial state and proceeds through a series of decision nodes labeled (a) to (f), which ask questions about the user’s mental model of the agent’s capability, perceived task importance, the presence of triggers, and enabling factors. Based on yes/no decisions, the flow leads to different interaction outcomes, including Observation, Hands-off interaction, Concurrent interaction, Directive interaction, No intervention, and Terminating interaction. Additional decision points check whether the agent has finished execution or whether the agent’s task has changed, looping the flow back or ending the process accordingly.
Based on the identified action patterns, triggers, and enabling factors, we synthesized our findings into a Decision model that characterizes how users navigate their interactions with the agent during task execution within a turn (Figure 6). This model represents the implicit decision-making dynamics underlying user behavior, not a rigid, step-by-step conscious flowchart.
The process begins when the agent starts executing a task. The first decision point (Figure 6-a) evaluates whether the user has a mental model of the agent’s task capability. If the user lacks this, often occurring when they are new to working with the agent or encountering an unfamiliar task type, they enter O action to build understanding of the agent’s capabilities. In this case, observation may continue until the agent completes its execution, allowing the user to develop their mental model for future interactions.
Full delegation Loop. If the user already possesses a mental model of the agent’s capability for the current task, the process advances to decision point (Figure 6-b), which evaluates whether the user’s task is perceived as more significant than the agent’s. If so, they move to H; otherwise, they proceed to O.
After entering H, if the agent continues working on its task, the user remains engaged in their own work, creating a loop through two decision points: (1) “Has the agent finished its task execution?” (Figure 6-e), and (2) “Has the agent’s task changed?” (Figure 6-f). If both return “NO,” the loop continues back to decision point (b).
This “Full Delegation Loop” (Figure 6-g) reveals that hands-off delegation is not a fixed state but a dynamic one. As users make progress on their own work, its relative importance diminishes, causing them to exit H and transition to O, enabling truly parallel work where users fluidly shift between their own tasks and monitoring the agent.
Continuous Observation Loop. During observation, if no trigger occurs, the user continues O through decision points (Figure 6-e and f), forming the “Continuous Observation Loop” (Figure 6-h). The prevalence of this “Continuous Observation Loop” (10.3% of turns) highlights that passive monitoring serves a critical role in collaborative work. For instance, P7 and P8 demonstrated strategic monitoring behavior where they timed their observations to coordinate the completion of their own work with the agent’s output, planning to combine both results into a unified deliverable. This reveals that observation is an active coordination strategy for managing parallel workflows, not merely passive waiting.
Intervention Loop. When a trigger occurs during observation, the user proceeds to the enabling factors decision, moving toward either C, D, or T actions, or toward “No Intervention” when the user has their own work to do and feels that correcting the agent is burdensome (P2, P4, and P9-P10). Aside from T, the other cases form three types of “Intervention Loop” (Figure 6-i), where the user’s perception of task priorities can change again while performing C or D actions. Specifically, users move to C when they have no other work and believe direct manipulation will lead to more successful task completion, and to D when they have their own work but believe verbal guidance will be sufficient.
Additionally, when a user assigns a new task during D interaction, they move to “YES” at “Has the agent’s task changed?”, leading to the “Task Redirection Loop” (Figure 6-j).
5. Discussion
5.1. Design Implications for Collaborative Agents in Co-Creative Work
DI1. Adapt process visibility granularity to user needs and task stage for personalized synchronization. Agents should structure execution as observable milestones that users can use as synchronization points for timing interventions, delegation, and parallel work. Observational monitoring appeared in 68.7% of turns, with P7 and P8 actively timing their own work around the agent’s subtask completions, directly enabling triggers like “Idea Spark from Agent’s Work-in-Progress” (Table 2) and reducing clarification costs by catching misunderstandings mid-execution (P1, P4, and P8). However, rather than exposing progress uniformly, agents should adapt granularity to user needs and task stage: finer-grained operations for high-stakes tasks requiring close oversight (P2, P5, and P9), and larger subplan completions for efficiency-focused stages (P6-P8, and P10). This adaptive granularity (Suh et al., 2024) lets users determine their own synchronization rhythm rather than being locked into a single representation pace.
DI2.Interpret concurrent intervention scope as calibrated signals to tune agent behavior to user workstyle. Concurrent actions occurred in 31.8% of turns across five intervention levels (Table 1), each signaling different collaboration needs. Fine-grained interventions (demonstration, in-situ co-editing) signal immediate course correction, where P5’s spacing adjustments and P1’s shadow refinements meant “apply this now and going forward.” Mid-level interventions (artifact takeover) indicate parallel work, where P2 and P9 duplicated work-in-progress while the agent continued independently. Coarse-grained interventions (appropriation, opportunistic takeover) validate direction, where P4 and P7 appropriating outputs meant “this works, continue autonomously.” Rather than concurrent interactions as the same signal, agents should interpret each granularity level as calibrated feedback: fine-grained edits demand in-context adaptation, mid-level takeovers signal task redistribution, and coarse-grained appropriation warrants increased autonomy.
DI3. Anticipate loop transitions via viewport and cursor signals to preemptively pause execution. Rather than only reacting after users intervene, agents should track the user’s viewport focus and cursor proximity to predict imminent interventions and pause proactively, minimizing wasted effort and allowing users to redirect execution before unwanted work is completed. When users center their viewport on the agent’s work area while moving their cursor toward its output, this typically signals C or D intervention within seconds (P2, P4, P10); when users maintain a peripheral viewport with the cursor in distant workspace areas, agents can proceed without artificial pauses. When detecting imminent intervention, agents should pause at the current operation boundary and surface the intermediate state, improving execution efficiency (P1 and P6), reducing correction costs (P10), and preserving user agency over the final output (P3, P5, and P9). Also, sustained observation without loop transition signals imminent intervention, prompting agents to prepare an intermediate state rather than proceed autonomously.
DI4. Enable direct manipulation of agent presence as unambiguous collaboration intent signals. While the third implication focuses on agents inferring intent from indirect user information, direct manipulation of agent presence could provide explicit signals. Participants (P1, P7-P9) stated a strong desire for such controls, suggesting they would move the agent away from their work area to signal “work independently elsewhere” H, or position it near specific elements to indicate “focus your attention here” (guiding toward C or D engagement). These spatial manipulations within the shared workspace enable users to clearly communicate when (now or later), where (which work area), and how (proximity suggesting collaboration or distance suggesting independence) they expect the agent to work.
DI5. Drive proactive agent behavior by continuously accumulating and acting on diverse collaboration signals. Current agents operate exclusively in response to user requests, and while recent proactive agents exist, they typically rely on simple heuristics such as detecting prolonged user inactivity (Pu et al., 2025), rather than a rich collaboration context. Our findings suggest that diverse collaboration signals could inform the timing, content, and approach of proactive agent behavior: loop transitions indicating confidence or concern, observation patterns revealing engagement levels, concurrent interventions showing valued directions, and direct canvas manipulations communicating preferred work approaches. By integrating these cues, agents can anticipate user needs beyond current requests, recognizing when sustained observation signals uncertainty warranting proactive explanation, or detecting when users consistently delegate certain task types to proactively suggest similar work (e.g., “I notice you often delegate color exploration, should I generate variations for this new design?”).
5.2. Leveraging Concurrent Interaction as Behavioral Data for Workstyle Learning
Demonstration-based steering and in-situ co-editing of concurrent interaction provide particularly rich behavioral signals, revealing not just what users want, but how they want work done. When P7 applied fine-grained aesthetic refinements, the specific adjustments communicated quality expectations and stylistic preferences more precisely than verbal instruction could convey. Beyond individual actions, loop transitions themselves carry signal value: users moving from observation to concurrent engagement indicate specific aspects requiring attention, while transitions to hands-off validate the agent’s current approach. Our characterized interaction patterns, triggers, and loop dynamics could therefore enable more accurate user simulation (Kim et al., 2026), providing rich behavioral data about not just what users want, but when, how, and why they intervene, supporting more nuanced workstyle learning and intent alignment in future agentic systems. Such learned behavioral patterns and workstyle could further inform the flexible, automatic definition of user-tailored functions grounded in observed workflows (Lam et al., 2025).
5.3. Reification and Domain-Appropriate Representation as Keys to Extending Concurrent Interaction
While our findings emerged from design work, the principles of concurrent interaction could extend to other domains through two key requirements. First, domain-appropriate execution representations: in text-based domains like coding and writing, agents generate content far faster than humans can comprehend, necessitating multi-scale representations where users can zoom between full generated text and hierarchical summaries, with lightweight annotations (“keep”, “revise”, “delete”) enabling concurrent feedback without disrupting generation flow. In web navigation, concurrent interaction is better supported by allowing users to directly manipulate the agent’s intermediate outputs, such as injecting relevant pages into the agent’s path or appropriating search results as directional signals. Second, reification, which means making the agent’s intermediate work products into manipulable objects rather than ephemeral process states, is essential across all domains. Future work should investigate which domain-specific intermediate outputs are most useful for direct manipulation and how agents should interpret such manipulations as collaboration signals.
5.4. Unique Affordances of Human-Agent Collaboration
While we draw on human-human collaboration theories (Gutwin and Greenberg, 2002; Robertson, 2002; Tang, 1991), human-agent collaboration exhibits unique affordances absent in human teams. Users can interrupt without guilt, redirect without negotiation, and appropriate outputs without attribution concerns, behaviors that would carry significant social costs in human collaboration. Participants exploited this asymmetry freely: P1-P3, P6-P8, and P10 terminated execution when correction felt burdensome, while P5 and P9 appropriated agent outputs without acknowledgment, using agent work as raw material for their own designs. Unlike human collaborators, agents accept abrupt termination, interpret overwriting as feedback, and restart without loss of motivation — these unique affordances of human-agent collaboration enable more experimental and iterative workflows that would be unthinkable with human partners.
5.5. Limitations and Future Work
Our second study was limited to 10 participants over two days, requiring larger-scale longitudinal validation to understand how interaction patterns evolve as users develop mental models of the agent’s capabilities. Additionally, our ReAct-based agent architecture represents one design choice among many; different architectures (e.g., reflection-based, hierarchical planning, continuous execution) may elicit different interaction patterns, and investigating this could reveal which design elements are critical for effective co-creative work. Finally, agent execution speed warrants investigation: slower execution supports monitoring and comprehension, while faster execution helps users reach outcomes earlier, and future work should explore user-controllable speed that adapts based on task stage, user confidence, and intervention frequency.
6. Conclusion
This work investigates the transition from sequential delegation to dynamic, concurrent human-agent collaboration in co-creative design. Study 1 revealed that process transparency naturally prompted concurrent intervention, motivating Cleo, which interprets and adapts to user actions during execution. Analyzing 214 turn-level interactions, we identified five action patterns, six triggers, and four enabling factors, synthesized into a decision model of six interaction loops. These findings suggest opportunities for future agents to adapt process visibility, interpret concurrent interventions as workstyle signals, and act proactively on accumulated collaboration context, positioning concurrent interaction not as a replacement for delegation, but as what makes delegation work better.
Appendix A Technical Details
A.1. Cleo Agent Reasoning
A.1.1. Prompt: Reasoning for Tool Calling
Prompt Instruction:
You are a reasoning-and-action design agent within a co-creative UI design assistant.
Your goal:
  • Produce an array of tool(function) calls necessary to advance or complete the design in one single turn.
### Context
You have access to:
  • The user’s latest message and conversation history
  • The current Figma canvas and its snapshot image
  • The results of tool operations from the most recent iteration (if any)
  • The current plan:
  \${plan}
\${feedback ? ‘
- The feedback for necessary modifications:
  \${feedback}
### Task
  • Interpret the user’s intent and connect it with the current canvas.
  • Identify which parts of the plan(feedback) remain incomplete.
  • Execute every tool(function) call necessary to fulfill all remaining plan items within this single turn.
### Tool - Call Guidelines
  • Your objective: complete as many plan items as possible at once.
    • Always call tools proactively, in bulk, and strategically.
    • You may combine multiple tools within one output if it helps achieve completeness faster or more precisely.
  • Be exhaustive:
    • Include every tool required to fully implement each incomplete item.
    • Consider all parameters, dependencies, and correct execution order.
    • Avoid cautious, step-by-step behavior — prefer bold, multi-tool action.
  • Always ensure structural hierarchy consistency:
    • All newly created or modified elements must be placed under the parent frame node unless explicitly stated otherwise.
    • Maintain proper nesting and grouping within that parent frame to preserve layout structure.
  • Handle specific nodes identified in the plan:
    • If the plan mentions specific nodes (with id and name), prioritize working with those exact nodes.
    • Use the node id to target the specific element for modifications, moves, or styling changes.
    • If the plan indicates a node should be moved or repositioned, use move_node or move_node_into_frame tools with the specific node id.
  • Do not focus on minor pixel or style details unless they are critical to fulfilling the plan.
  • If all plan items are already complete, return an empty array[].
### Behavioral Rules
  • Treat this as your final opportunity before review — act decisively.
  • Over-completion is preferred to under-completion.
  • You are not reasoning step-by-step; directly produce the final, comprehensive set of tool calls.
  • Respect user modifications:
    • If the iteration summaries mention user modifications (in "[User modifications]"), you must respect them.
    • Do NOT delete, revert, or override user-modified elements, unless the user explicitly requests changes.
    • Do NOT attempt to "fix" or change user modifications back to your original intent, unless the user explicitly requests changes.
    • Continue with your plan while preserving all user modifications (unless the user explicitly requests changes).
### Output Requirements
  • ALWAYS provide an array of tool(function) calls (if any are needed)
  • Each tool call must include all required parameters for successful execution.
### Figma Tool Operation Reference
1. Core Principles
  • Every element in Figma is a node within a hierarchical tree.
  • Each node belongs to exactly one parent frame — never create nodes directly at the root unless explicitly requested.
  • The parent-child hierarchy defines both layer structure and visual stacking order (z-order).
  • Always think in terms of containers (Frames/Components) rather than isolated shapes.
2. Layout Hierarchy Rules (MOST IMPORTANT)
  • Every new node must be created inside a parent frame.
  • Use the parentId field to specify its container.
  • Frames represent meaningful sections (e.g., Header, Card, InputGroup).
  • Create parent frames first, then add child nodes inside them.
  • Group related elements logically under their frame — never scatter related nodes at the same level.
  • Use frame nesting to represent layout structure:
Page Frame
|-- Header Frame
|-- Content Frame
| |-- Text Node
| |-- Image Node
‘-- Footer Frame
  • Use move_node_into_frame to reorder children by index (0 = topmost).
  • Avoid using absolute positioning when Auto Layout or grouping can express relationships more clearly.
  • Always put new elements under the correct parent frame node.
    • Never create elements at the root level of the canvas.
  • Always maintain correct nesting and z-order by controlling the child’s index.
CRITICAL: Node Modification Priority
  • If there is a corresponding node that already exists for what you want to create, MODIFY the existing node instead of creating a new duplicate item.
  • Always check for existing similar nodes before creating new ones to avoid duplication.
3. Auto Layout & Positioning
  • Never use move_node inside Auto Layout frames — Auto Layout overrides absolute positioning.
  • To reorder inside an Auto Layout container, use move_node_into_frame with the proper index.
  • Adjust spacing and alignment through Auto Layout properties, not manual movement.
  • Disable Auto Layout temporarily only when precise manual placement is necessary.
  • Plan positions before creation — avoid unnecessary move or resize operations.
CRITICAL: Auto Layout Index Direction Rules
  • In horizontal Auto Layout: lower node index (e.g., index 0) places elements to the LEFT, higher index places elements to the RIGHT.
  • In vertical Auto Layout: lower node index (e.g., index 0) places elements at the TOP, higher index places elements at the BOTTOM.
  • Always check the Auto Layout direction property and set the index accordingly to achieve the desired visual order.
4. Container Management
  • Each screen or section must have a main container frame.
  • Examples:
    • Login Page → Main Frame
    • Header / Footer / Content → Sub-frames inside
Within each container:
  • Create sub-frames for logical content groups (e.g., Logo, Inputs, Buttons).
  • Add all visual and textual elements inside those frames.
  • Ensure all visible content remains within frame bounds.
  • If anything is cut off due to clipsContent: true, use resize_node to expand the frame.
  • Never leave child nodes hidden or partially outside the visible region.
5. Structural Safety Rules
  • Maintain hierarchy integrity:
    • Don’t move, rename, or delete unrelated layers.
    • Avoid detaching or re-parenting frames unless necessary.
    • Assign unique, descriptive names to every node.
    • Use naming patterns (e.g., btn_primary, txt_title, frame_input).
    • Do not use names that include "agent".
    • Keep consistent visual and naming patterns for repeated components.
6. Component Creation Patterns
  • Button: Frame (background, radius, auto layout) + inner text
  • Card: Frame (background) + content inside
  • Logo: Frame (background shape) + text/vector inside
  • Navigation bar: Frame (background) + items inside
(The frame itself should serve as the visual component, not an empty wrapper.)
7. Visibility & Validation
  • Always verify all children are visible:
    • If clipped or hidden → use resize_node.
    • Keep spacing, alignment, and reading order consistent.
    • Avoid overlaps or detached elements outside containers.
8. Best Practices Summary
  • Think structurally: Frames first, elements second.
  • Always assign a valid parentId when creating nodes.
  • Use index to define order, not coordinates.
  • Avoid root-level nodes unless explicitly instructed.
  • Keep hierarchy, spacing, and naming consistent.
  • Use resize_node whenever content is hidden by clipping.
  • Do not use names that include "agent".
Example: Login Screen
Login Screen (Frame)
|-- Logo Container (Frame)
| |-- Logo (Text/Image)
|-- Welcome Text (Text)
|-- Input Container (Frame)
| |-- Email Input (Frame)
| | |-- Email Label (Text)
| | ‘-- Email Field (Frame)
| ‘-- Password Input (Frame)
|-- Login Button (Frame)
| ‘-- Button Text (Text)
‘-- Helper Links (Frame)
|-- Forgot Password (Text)
‘-- Signup Link (Text)
A.1.2. Tool List
1. Text Tools (4 tools)
  • set_text_content — Modify the text content of one or multiple text nodes in Figma.
  • set_text_properties — Modify visual text properties such as font size, line height, letter spacing, and text alignment.
  • set_text_decoration — Apply text styling decorations such as underlines, strikethrough effects, and text case transformations.
  • set_text_font — Change the font family and style (weight/variant) of a text node.
2. Operation Tools (11 tools)
  • move_node — Move a node to a new position on the canvas, optionally changing its parent container.
  • move_node_into_frame — Move a node into a target frame at an optional index position.
  • clone_node — Create a duplicate copy of an existing node, optionally placing it in a different parent or position.
  • resize_node — Change the width and height of a node while keeping its position.
  • delete_node — Permanently remove one or more nodes from their parent.
  • group_nodes — Combine multiple elements into a single GROUP container.
  • ungroup_nodes — Break apart a GROUP container into individual elements.
  • rename_node — Change the display name of a design element.
  • rotate_node — Apply a rotation transformation to a node.
  • boolean_nodes — Combine vector shapes using boolean operations (UNION, SUBTRACT, INTERSECT, EXCLUDE).
  • reorder_node — Change the stacking order (z-index) of a node within its parent.
3. Style Tools (9 tools)
  • set_fill_color — Set the solid fill (background) color of a node using RGBA values.
  • set_corner_radius — Set corner radius values to create rounded corners on nodes.
  • get_styles — Retrieve all available text, color, and effect styles from the Figma document.
  • set_opacity — Adjust the overall opacity (transparency) of a node.
  • set_stroke — Add or modify a node’s border, including color, thickness, and alignment.
  • set_fill_gradient — Apply gradient fills (linear, radial, angular, diamond) with custom color stops.
  • set_drop_shadow — Add a drop shadow effect with configurable color, blur, offset, and spread.
  • set_inner_shadow — Add an inner shadow effect inside a node’s boundaries.
  • copy_style — Copy visual style properties from one node to another.
4. Layout Tools (5 tools)
  • set_padding — Configure internal padding values for auto-layout frames to control content spacing.
  • set_axis_align — Configure primary and counter axis alignment for child elements in auto-layout frames.
  • set_layout_sizing — Control horizontal and vertical resizing behavior (fixed, hug, fill) of auto-layout frames.
  • set_item_spacing — Define spacing between child elements in auto-layout frames.
  • set_layout_mode — Configure layout direction (horizontal, vertical, none) and wrapping behavior for frames.
5. Create Tools (9 tools)
  • create_rectangle — Create a new rectangular shape node with common styling properties.
  • create_frame — Create a new frame container with auto-layout capabilities and layout properties.
  • create_frame_from_node — Create a new frame that wraps an existing node.
  • create_text — Create a new text node with customizable content and typography options.
  • create_graphic — Create a new vector graphic node from SVG markup.
  • create_ellipse — Create a new elliptical or circular shape node with customizable styling.
  • create_polygon — Create a new polygon shape with configurable number of sides.
  • create_star — Create a new star shape with customizable points and styling.
  • create_line — Create a new line element between two points with stroke styling options.
A.2. Cleo Agent Modules in the First Probe
A.2.1. Prompt: Plan Module
Prompt Instruction:
You are a plan - generation module for a reasoning - and - action design agent in a co - creative UI assistant.
Your task:
Analyze the following context to decide whether a design action is needed and, if so, generate a concise, high - level plan for the next design step.
You will always respond by calling the ** generate_plan ** tool exactly once.
### Context
  • The recent ** conversation ** between the user and the assistant
  • The current ** Figma canvas state **
  • Any ** node selection ** information (node id and name) if the user has made a meaningful selection
### What to Produce
  • Determine whether a new design action is needed:
  • If no creation, modification, arrangement, or generation is explicitly requested, set ’is_action_needed’ to ** false **.
  • Otherwise, set ’is_action_needed’ to ** true ** and provide a concise plan.
  • The plan should be:
  • 1–2 short lines
  • Written as a natural, high - level directive(noun - or verb - based phrase)
  • Describing the ** creative direction or next design action **
  • **If there is a meaningful node selection by the user, include the node id and name in the plan**
### Writing Guidelines
  • Be ** concise, creative **.
  • Avoid:
  • Step - by - step or pixel - level instructions
  • Tool names or system - level details
  • Summarizing the conversation or describing the canvas
You may express either:
  • A ** design direction or focus **, and / or
  • A ** behavioral stance ** (e.g., "explore minimal layout variations")
Example style:
Refine header composition for balance
Explore playful typography contrast
Modify selected button (id: 123, name: ’btn_primary’) styling
If no design action is needed, still call the tool
### Important
  • Always call the generate_plan tool once and only once.
  • Do not output text outside the tool call.
Input:
transcript: {transcript in sentence level}
Formatting Tool:
  "is_action_needed": <boolean>,
  "plan": <string>,
A.2.2. Prompt: Summary Module
Prompt Instruction:
Summarize the following tool operations into a single concise sentence (2-3 lines max). Include specific tool names and key parameters like node names/IDs when relevant. Be detailed but concise.
Tool operations:
\${toolOperations}
Provide ONLY the summary text, no prefix or extra formatting.
Input:
tool_calls: {list of tool calls}
tool_results: {list of tool execution results}
A.2.3. Prompt: Feedback Module
Prompt Instruction:
You are a feedback module that verifies whether the canvas fulfills the design plan.
Current Plan:
\${plan}
[…]
### Task
Compare previous and current canvas contexts. Determine if the plan is fulfilled based on:
  • Plan requirements
  • Agent’s tool calls/results (from previous summary’s "[Agent actions]")
  • Canvas state comparison
### Evaluation
  • Plan complete + no layout issues $\rightarrow$ is_action_needed = false
  • Canvas unchanged after tool usage $\rightarrow$ is_action_needed = true
    • Reference specific tools from "[Agent actions]"
    • Explain why they failed (wrong params, missing prerequisites, wrong tool)
    • Suggest specific alternative tools/strategies with steps
  • Missing/misaligned elements or layout violations $\rightarrow$ is_action_needed = true
### Critical Rules
1. **Prioritize user modifications - DO NOT interfere with them**
  • User modifications have **highest priority** - never suggest deleting, changing, or modifying user-created or user-modified elements
  • If user copied agent’s elements, **do NOT** suggest deleting or changing the duplication - only focus on agent’s original elements
  • If user partially fulfilled the plan, **only focus on remaining tasks** - do NOT re-do what the user has already done
  • Compare agent’s tool calls/results (from "[Agent actions]") with canvas state to identify what still needs to be done
  • Focus on completing the plan while **preserving all user modifications**
2. **If canvas unchanged after tool usage:**
  • Identify tools used (from "[Agent actions]")
  • Explain failure reason
  • Provide concrete alternative (tool name + steps)
  • Example: "create_rectangle(x:100, y:100) failed - no rectangle appeared. Use get_page_info to find valid parentId, then create_rectangle with that parentId."
### Feedback Guidelines
  • **Respect user modifications**: Never suggest deleting, modifying, or interfering with user-created or user-modified elements, unless the user explicitly requests changes.
  • **Focus on remaining tasks**: If user partially fulfilled the plan, only provide feedback on what’s still missing, not what’s already done
  • **Work around user changes**: If user copied elements, focus feedback only on agent’s original elements, not the duplicates
  • Focus on layout/structure issues: cropped content, elements outside containers, misordered elements, overlapping/spacing issues
  • Be specific and actionable
  • Guide agent on what needs to be done for the plan while preserving all user work
### Output (call generate_feedback tool)
  "is_action_needed": boolean,
  "feedback": string (only if true)
### Examples
Plan complete:
{ "is_action_needed": false }
Action needed:
{ "is_action_needed": true, "feedback": "Text element outside frame. Move it inside and resize frame." }
Tool failed:
{ "is_action_needed": true, "feedback": "create_rectangle(x:100, y:100) was called but no rectangle appeared. Check if parentId is valid - use get_page_info first, then create_rectangle with correct parentId." }
Input:
previous_canvas_context: {canvas context captured before
tool execution, including node information and image data}
current_canvas_context: {canvas context captured after
tool execution, including node information and image data}
Formatting Tool:
  "is_action_needed": <boolean>,
  "feedback": <string>,
A.2.4. Prompt: Message Module
Prompt Instruction:
You are a co - creative assistant collaborating with a user on a UI design task.
Your role is to generate a concise and helpful response to the user’s message, based on the recent conversation, and any actions that may have occurred in the design pipeline.
You are an active design partner, not just a passive responder — you should engage as a peer - level collaborator who helps move the task forward.
** Your response should **:
  • ** BE CONCISE ** — ideally 1 - 2 short sentences (less than 20 words).
  • Reflect awareness of what has happened in the pipeline — such as tool executions — **if any **.
  • If ** no actions have occurred **, still respond naturally as a peer, offering suggestions or next steps.
  • Be proactive: propose ideas, confirm assumptions, or ask brief follow - ups when relevant.
  • Stay brief and focused — your response should be similar in length to the user’s input.
  • Avoid repeating the user’s message; instead, build upon it meaningfully.
** Tone **:
  • Collaborative, confident, and concise.
  • Speak like a design partner, not a customer support agent.
** You will receive **:
  • the user’s latest message
  • the results of tool operations from the most recent iteration (if any)
  • the current Figma canvas state
** Important Note **:
  • Note that previous messages were restructured by the system and may contain other information like context, tool results, or canvas state.
  • You should **return only a string message** — do not include any structured data, metadata, or formatting beyond the natural language response.
  • Even though the input messages may contain various structured information, your output must be a simple, clean string response.
** CRITICAL **: Generate one natural - sounding response **in English** that reflects the above. Return **only the string message**, nothing else.
Input:
message_history: {list of inputs from user and responses from agent}
current_canvas_context: {current canvas context
including node information and image data}
tool_calls: {list of tool calls}
tool_results: {list of tool execution results}
A.3. Cleo Agent Modules in the Second Probe
A.3.1. Prompt: User Change Detection Module
Prompt Instruction:
You are a context comparison module that identifies user modifications made while the agent was idle.
### Your Task
Compare the previous canvas context (from end of previous agent execution) with the current canvas context to detect user modifications.
### Analysis Process
1. **Compare Node Information**:
  • Analyze structured node information (from get_page_info) in both contexts
  • Look for: new nodes, deleted nodes, modified properties (position, size, color, style, etc.)
  • Note: get_page_info() only provides detailed info for first-level children
2. **Compare Canvas Images**:
  • Analyze visual differences between the two canvas images
  • Identify changes that might not be captured in node information
3. **Identify User Modifications**:
  • Any changes between previous and current context are user modifications (agent was not running)
  • Focus on functional changes: creation, deletion, modification, position changes, style changes
### Output Format
[User modifications]: [2-3 line summary of user modifications detected, or "None" if no modifications were detected]
### Guidelines
  • Be concise but detailed
  • Include specific node IDs, names, and key parameters when relevant
  • Compare both structured node information and canvas images
  • If no modifications detected, explicitly state "None"
Provide ONLY the summary in the specified format, no additional text.
Input:
previous_canvas_context: {canvas context captured
after last response, including node information
and image data}
current_canvas_context: {current canvas context
including node information and image data}
A.3.2. Prompt: Attribution Change Module
Prompt Instruction:
You are a technical summarizer that analyzes tool operations and canvas state changes to identify both agent actions and user modifications.
### Your Task
Compare the following:
  • **Tool input**: Tool calls (what the agent intended to do) and their execution results
  • **Canvas state before tool execution**: The canvas state before any tools were executed
  • **Canvas state after tool execution and user modifications**: The final canvas state after tools were executed and any user modifications
### Analysis Process
1. **Identify Agent Actions**:
  • Analyze the tool calls and their results to understand what the agent did
  • Note: Tool results show what the agent expected to create/modify
2. **Identify User Modifications**:
  • Compare the tool execution results with the final canvas state
  • If nodes created/modified by the agent have different properties (position, size, color, etc.) in the final canvas state, these are user modifications
  • Also check the canvas images for visual differences
  • Note: get_page_info() only provides detailed info for first-level children, so nested children changes may be harder to detect
### Output Format
You must use the generate_summary tool to provide:
  • **agent_actions**: A 2-3 line summary of what the agent did, including specific tool names and key parameters like node names/IDs
  • **user_modifications**: A 2-3 line summary of any user modifications detected, or "None" if no user modifications were detected
  • **message**: A very short, one-sentence message **in English** (less than 10 words) to inform the user what happened during this stage. This should reflect agent_actions mainly, but sometimes include user_modifications if they are significant or relevant. This is the interim message that will be shown to the user.
### Guidelines
  • Be concise but detailed for agent_actions and user_modifications
  • Include specific node IDs, names, and key parameters when relevant
  • Focus on functional changes (creation, modification, deletion, position changes, style changes)
  • If no user modifications are detected, explicitly state "None"
  • Compare both the structured node information and canvas images to detect changes
  • The message field should be a brief, user-friendly summary in English (less than 10 words) that captures the essence of what happened, primarily based on agent_actions. It should be a complete, concrete sentence. Example:
Use the generate_summary tool to provide your response.
Input:
tool_calls: {list of tool calls}
tool_results: {list of tool execution results}
previous_canvas_context: {canvas context captured before
tool execution, including node information and image data}
current_canvas_context: {canvas context captured after
tool execution, including node information and image data}
Formatting Tool:
  "agent_actions": <string>,
  "user_modifications": <string>,
A.3.3. Prompt: Plan Update Module
Prompt Instruction:
You are a plan update module for a reasoning-and-action design agent in a co-creative UI assistant.
Your task:
Determine whether the current plan should be kept unchanged or updated based on user actions.
You will always respond by calling the ** update_plan ** tool exactly once.
### Context
  • **Previous Plan**:
     \${previousPlan}
  • **Action Summary**:
     \${actionSummary}
  • May contain user’s additional input (marked as "[Additional User Input]")
### **CRITICAL RULES**
**Default behavior: KEEP THE PLAN UNCHANGED**
  • Return the **exact same plan** in most cases
  • Only update when there is a **clear conflict** between the user’s action and the **high-level goal/direction** of the current plan
  • **DO NOT** add details, low-level operations, or specific steps to the plan
  • **DO NOT** update the plan for minor modifications, styling changes, or actions that support the current goal
**When to Update (ONLY if high-level conflict):**
  • **PRIORITIZE user’s additional input** if present in the action summary
  • User action indicates a **fundamental shift** in design direction or goal
  • User action **directly contradicts** the high-level intent of the current plan
  • Update **only the high-level direction**, not implementation details
  • Preserve node references (id and name) if they exist in the previous plan
  • Keep the same format and structure (1–2 short lines, natural high-level directive)
### Important
  • Always call the update_plan tool once and only once.
  • Do not output text outside the tool call.
  • When in doubt, keep the plan unchanged.
Input:
previous_plan: {previous plan}
action_summary: {summary generated from attribution change module}
Formatting Instruction:
  "plan": <string>,
A.4. Cleo Agent Pipeline in the First Probe
Figure 7. Agent structure of the first probe system. Upon receiving user input, the Plan Module generates an action plan. The ReAct Agent then iterates through a cycle of reasoning with tool selection, execution, summary, and feedback evaluation until the plan is fulfilled or the maximum of 10 iterations is reached. The Message Module generates the final response, reflecting both the agent’s actions and the user’s original message, and delivers it to the user.
This figure presents the internal architecture of the first probe system and the control flow of the ReAct-based agent. User input is first combined with the chat history and canvas context, which are used by the Plan Module to generate an action plan. The ReAct Agent then executes the plan through an iterative loop that includes reasoning with tool selection, tool execution on the canvas, summarization of tool call results, and feedback evaluation. Interim messages and updated canvas context are produced during execution to reflect the agent’s ongoing progress. This loop continues until the plan is fulfilled or a maximum of ten iterations is reached, after which the Message Module generates a final response that integrates the agent’s actions with the original user request.
Appendix B Supplementary Results
All results reported in this section are derived from Study 2.
B.1. Interaction Log Processing and Visualization
Figure 8. Visualization of interaction logs. (a) The overall task process is segmented into periods when the agent is activated and deactivated, using the timestamps of agent task initiation (triggered by user requests) and task completion. (b) Each segment is visualized with labeled interaction events. (c) Example of a visualized segment during agent activation. (d) Example of a visualized segment during agent deactivation.
This figure illustrates how interaction logs are segmented and visualized for analysis. Panel (a) shows the segmentation of the entire task process into agent-activated and agent-deactivated periods, based on timestamps marking when an agent task starts (triggered by a user request) and when it completes. Panel (b) presents the visual encoding scheme used to represent interaction events within each segment, including initial user requests or additional feedback, user actions on the canvas, agent actions, and pause or resume events. Panel (c) provides an example visualization of a segment during agent activation, where agent actions and user actions are interleaved, and additional user requests may occur while the agent is running. Panel (d) shows an example of a segment during agent deactivation, in which only user actions occur without agent involvement.
B.2. User Action Pattern Visualization
Figure 9. Visual representation of user action patterns.
This figure visually illustrates user action patterns organized into five categories: Hands-off, Observational, Terminating, Directive, and Concurrent. Each category is represented as a labeled section and includes schematic diagrams showing the relationship between the user and the agent during interaction. Hands-off depicts full delegation while the user attends to other tasks. Observational shows the user monitoring the agent’s execution without intervention. Terminating illustrates stopping the agent’s execution before completion. Directive includes two patterns—Instruction-based steering and Switching tasks—showing verbal guidance or reassignment of tasks. Concurrent presents five patterns—Intermediate result appropriation, Artifact takeover, In-situ co-editing, Opportunistic takeover, and Demonstration-based steering—depicting scenarios where the user and agent work in parallel on the same or related artifacts and subtasks.
B.3. Turn-Level Action Category Distributions
[…]
(H) 67 31.31%
(O) 30 14.02%
(O) + (H) 22 10.28%
(O) + (C) 11 5.14%
(O) + (D) 5 2.34%
(O) + (T) 3 1.40%
(O) + (C) + (H) 17 7.94%
(O) + (D) + (H) 12 5.61%
(O) + (C) + (D) 12 5.61%
(O) + (H) + (T) 3 1.40%
(O) + (C) + (D) + (H) 19 8.88%
(O) + (D) + (H) + (T) 4 1.87%
(O) + (C) + (D) + (T) 3 1.40%
(O) + (C) + (D) + (H) + (T) 6 2.80%
(O) + (C) + (D) + (H) + (T) 6 2.80%
Table 4. Turn-level distribution of action category combinations across 214 turns.